### v1

# rtboxstyle linecolor=black, backgroundcolor=lightlight-gray, leftmargin=0cm, rightmargin=0cm, skipabove=4pt, skipbelow=4pt, linewidth=0.4pt, innertopmargin=3pt, innerbottommargin=3pt, innerrightmargin=6pt, innerleftmargin=6pt, roundcorner=2

Abstract
Provenance graph analysis plays a vital role in intrusion detection, particularly against Advanced Persistent Threats (APTs), by exposing complex attack patterns. While recent systems combine graph neural networks (GNNs) with natural language processing (NLP) to capture structural and semantic features, their effectiveness is limited by class imbalance in real-world data. To address this, we introduce ProvSyn, an automated framework that synthesizes provenance graphs through a three-phase pipeline: (1) heterogeneous graph structure synthesis with structural-semantic modeling, (2) rule-based topological refinement, and (3) context-aware textual attribute synthesis using large language models (LLMs). ProvSyn includes a comprehensive evaluation framework that integrates structural, textual, temporal, and embedding-based metrics, along with a semantic validation mechanism to assess the correctness of generated attack patterns and system behaviors. To demonstrate practical utility, we use the synthetic graphs to augment training datasets for downstream APT detection models. Experimental results show that ProvSyn produces high-fidelity graphs and improves detection performance through effective data augmentation.
1 Introduction
Data provenance [29], which involves recording the history of data modifications, access, and usage, serves as a critical tool in cybersecurity systems for tracing attacks [61]. However, as system scales expand, the volume of provenance data increases dramatically, posing challenges for data management and analysis. To address this, researchers have begun to represent the provenance of the data using graph structures and applying graph machine learning algorithms for the detection and tracking of intrusions [21, 62, 53]. This approach effectively captures the complexity of system activities and enables the detection of attacks that traditional methods struggle to identify, such as zero-day attacks and APT attacks.
The quality of provenance data sets largely determines the effectiveness of provenance-based intrusion detection systems (IDS) trained on them. However, there are very few datasets available, and even fewer datasets that can accurately describe benign and malicious activities in a system. One possible solution is to produce system logs captured from real-world scenarios. However, one of the key disadvantages of using such datasets is privacy concerns. Sharing such dataset across the community could result in leakage of critical privacy information, and hence it’s difficult for organizations to release such datasets. Another solution is to collect benign and malicious activities in experimental environments. For example, the DARPA Transparent Computing (TC) [2] program contributes two widely used datasets in the cybersecurity community, DARPA TC dataset and DARPA OpTC dataset. However, such datasets, while giving us insights into system activities, convey limitations, one of which being the class imbalance issue [6], where less common or underrepresented program behaviors are under-sampled. This imbalance hampers the generalization ability of machine learning models trained on such data, limiting their effectiveness in unseen environments [77, 9, 76, 6].
To mitigate the issue of data imbalance, a common approach is to synthesize high-fidelity provenance data that expands coverage of minority classes and patterns. Such data synthesis can be achieved through methods such as graph generation models or large language models (LLMs), yet each has critical limitations. On the one hand, most graph generation models focus on structural connectivity and lack support for generating rich textual attributes, which are essential for downstream intrusion detection tasks [12, 71, 37, 53]. On the other hand, although LLMs are strong at producing semantically rich log entries, their outputs often fail to preserve the underlying graph structure when transformed into provenance graphs [4, 1].
To overcome these challenges, we propose ProvSyn, a hybrid framework that integrates the structural modeling ability of graph generation models with the text generation ability of LLMs to jointly synthesize provenance graphs with both accurate topology and meaningful textual attributes. Specifically, ProvSyn is a three-phase synthesis framework comprising structural synthesis, topological refinement, and textual attribution. The initial phase constructs heterogeneous graph topology by generating nodes and edges with categorical attributes. Subsequent structural refinement applies domain-specific rules to eliminate unreasonable connections and isolated entities. In the third phase, an LLM is used to generate textual attributes of nodes. However, the implementation of our framework confronts three principal technical challenges:
  • Scalability to provenance graphs. Provenance graphs are heterogeneous graphs with more than 10,000 nodes and edges, whereas most heterogeneous graph generation models are designed for molecular graphs and are typically trained on graphs with fewer than 100 nodes [72, 17]. Effectively scaling graph generation models designed for other domains to large-scale provenance graphs thus remains a challenge.
  • Textual attributes synthesis. First, LLMs inherently struggle with graph structure comprehension [33], necessitating transformation of provenance graphs into LLM-interpretable formats. Second, due to lacking prior knowledge of typical provenance entity names, direct LLM prompting produces simplistic, low-diversity outputs.
  • Evaluation framework gap. The absence of a multidimensional evaluation framework hinders comprehensive assessment of synthesized graphs, particularly in verifying the semantic correctness of attack patterns and system behavior within generated provenance structures.
To address the first challenge, we adopt a domain-agnostic heterogeneous graph generation network [17] to jointly model node and edge labels along with structural connectivity. To make the training process scalable to large graphs, we adopt a restart-based random walk for subgraph sampling. Through a novel restart mechanism, this sampling algorithm escapes local regions and captures subgraphs that better reflect the overall graph structure.
To address the second challenge, we convert the graph structure into depth-first search (DFS) sequences to enhance the LLM’s understanding of the graph. In addition, we also designed two different masking strategies, including fully masking and partially masking, to construct a training set for fine-tuning the LLM, thereby providing it with knowledge of entity names in the provenance graph.
To address the third challenge, we design a comprehensive framework that evaluates fidelity from five perspectives: structural, textual, temporal, embedding, and semantic correctness. For structural fidelity, we adopt five Mean Discrepancy (MMD)-based metrics. For textual fidelity, we use standard NLP metrics, including BLEU, GLEU, and ROUGE. For temporal fidelity, we use sequence comparison metrics including Longest Common Subsequence (LCS) and Dynamic Time Warping (DTW). For embedding fidelity, we utilize various embedding methods, including DeepWalk and Doc2Vec. To accurately assess semantic correctness—the most challenging aspect—we adopt a model-based approach. Specifically, we sample positive and negative triples from real provenance graphs and train a classification model using contrastive learning.
In the experiments, we used our proposed evaluation framework to evaluate the provenance data synthesized by LLMs, including GPT-4o and Claude 3.7, as well as by ProvSyn. The results show that ProvSyn achieves higher fidelity than these strong LLMs. Moreover, by generating 100% novel and 100% diverse graphs, ProvSyn effectively mitigates the data imbalance issue in provenance datasets. When applied to an intrusion detection system, models trained on datasets augmented with ProvSyn-generated graphs exhibited reduced false positive rates during testing.
The main contributions of our work are as follows:
  • We propose ProvSyn, a three-phase provenance graph synthesis framework that enables the construction of high-fidelity heterogeneous graphs with rich textual attributes.
  • We design a comprehensive multi-dimensional evaluation framework that evaluates the fidelity of synthesized provenance graphs, including: structural, textual, temporal, embedding, semantic correctness.
  • We mitigate the data imbalance problem in provenance datasets by synthesizing novel and diverse provenance graphs, thereby enhancing the generalizability of downstream intrusion detection models.
  • The proposed generation and evaluation frameworks can also be extended to other domains involving Heterogeneous Information Networks (HINs), such as social networks and citation networks.
2 Background and Motivation
2.1 Class imbalance in Provenance Graphs
We statistically analyzed the distributions of entity type and event type across four commonly-used public provenance datasets (in Appendix A Figure 3). Among the datasets, Cadets, Theia, and Trace are collected from DARPA Engagement 3, and Nodlink [37] is simulated provenance data collected on an Ubuntu system using Sysdig. The distributions exhibit a typical long-tail pattern and reveal significant data imbalance. In both entity type and event type, a single category accounts for over 50% of instances, exceeding 70% in some cases. In contrast, most other categories contribute less than 20%, with many below 5%. This imbalance further reflects the uneven distribution of system behaviors and attack patterns in most provenance datasets.
The class imbalance problem poses a significant challenge for learning-based intrusion detection [21, 62, 74, 22, 39]. Previous studies in the deep learning community have revealed that class imbalance has a detrimental effect on a model’s classification performance as well as its generalization abilities [8, 34]. When deployed in open-world out-of-distribution (OOD) environments, detection models trained on imbalanced datasets may significantly exhibit performance degradation [67, 25]. Hence, a more balanced dataset is vital for deep learning-based provenance-based IDS (PIDS) to achieve optimal results. However, this problem is largely ignored by the community, as there are very few studies considering this issue.
2.2 Problems with Existing Methods
To mitigate class imbalance, conventional techniques such as undersampling [44], oversampling [47], and the Synthetic Minority Over-Sampling Technique (SMOTE)  [9] have been proposed. However, these methods are not suitable for provenance graph data. Undersampling breaks the causality chain between system entities, while oversampling does not add variety to benign activities and can cause overfitting. SMOTE, which relies on linear interpolation in the feature space, cannot be applied to provenance graphs due to overlapping distributions of different classes in the feature space.
Another approach for augmenting provenance datasets is to leverage graph generation models to synthesize provenance graphs. However, most existing models focus primarily on generating the structural connectivity among nodes and edges, without supporting additional node or edge attributes [12, 71]. In contrast, provenance graphs contain rich textual attributes associated with nodes and edges, and the text attributes are crucial features for downstream intrusion detection systems [37, 53]. Yet, to date, no graph generation model inherently supports the joint generation of both graph structure and textual attributes.
With the recent advances in large language models, their capabilities in natural language understanding and generation have been widely demonstrated [4, 1]. This motivates the exploration of the use of LLMs to synthesiz provenance data. Specifically, LLMs can be tasked with creating new provenance log entries. These generated logs contain rich texual and semantic information. However, when converting LLM-generated logs into graphs and evaluating them using graph-based metrics, their structural fidelity falls significantly short of graphs produced by heterogeneous graph generation models (see details in Section 5.2). This indicates that although current LLMs perform well in text synthesis, they lack a strong capability for synthesizing accurate graph topology and structure.
To address these limitations, we propose ProvSyn, a framework that integrates the complementary strengths of graph generation models for structural synthesis and large language models for text attribute generation.
3 Related Work
Provenance-based Intrusion Detection System Provenance-based detection methods can be divided into three main categories [74]: rule-based, statistics-based, and learning-based techniques. Rule-based methods [24, 46, 28] rely on predefined heuristics derived from known attack patterns to identify malicious activities. Statistics-based [23, 61, 45] approaches analyze deviations in provenance graph elements by measuring their statistical anomalies. Learning-based techniques employ deep learning models to capture either normal system behavior [21, 62] or malicious patterns [41, 74, 39], framing APT detection as either a classification [39] or anomaly detection task [74, 22]. These learning-based methods further branch into sequence-based approaches [41], which focus on execution workflows, and graph-based [21, 62, 74, 22, 39] methods, which use graph neural networks to model entity relationships and detect behavioral anomalies.
Network Traffic Augmentation To address class imbalance in network traffic datasets, one line of work applies sampling techniques. Jiang et al. [32] adopt one-side selection to remove noisy majority samples and apply SMOTE [9] to augment minority instances. Liu et al. [42] propose the Difficult Set Sampling Technique to improve sample representativeness. Another line of research explores synthetic network traffic generation using deep learning. Generative models, especially GANs, have been employed for various scenarios, including flow-based traffic [55], manipulated traffic [11], and social media traffic [54], addressing challenges in realistic traffic synthesis [5].
LLM-driven Synthetic Data Generation LLM-based synthetic data generation methods can be categorized into three main types: prompt engineering-based approaches [16, 20, 57, 43, 70, 38, 68], which utilize task descriptions, condition-value prompts, and in-context examples to steer generation; multi-step generation-based approaches [64, 15, 14, 63, 27], which break down complex tasks either at the sample or dataset level; and knowledge-enhanced approaches [50, 10, 30], which incorporate external knowledge such as knowledge graphs or web sources to improve the factual quality of outputs. The evaluation of synthetic data includes both direct metrics, which measure faithfulness [36] and diversity [73], and indirect metrics, which assess performance on downstream tasks through benchmark testing [58] and open-ended evaluation using human [26] or model-based [65] methods.
4 ProvSyn Design
We implement the generation of provenance graphs in three stages. In the first stage, we train a heterogeneous graph generation model using real provenance graph dataset to produce the graph structure. After that, we apply rule-based post-processing to remove invalid edges and isolated nodes. In the third stage, we fine-tune a large language model to label the nodes with appropriate names.
4.1 Problem Definition
A system audit log dataset can be represented as a provenance graph, where nodes correspond to system entities and directed edges represent system events. Formally, a provenance graph is a directed heterogeneous graph $G=(V,E)$, where each node $v_{i}\in V$ is associated with a type $t_{i}=f(v_{i})$ and a name $n_{i}=h(v_{i})$, and each edge $e_{i}\in E$ is associated with a type $l_{i}=g(e_{i})$. Each system event is represented as a labeled 5-tuple:
$e_{i}=(t_{i},n_{i},t_{j},n_{j},l_{i})$
where $t_{i}$ and $t_{j}$ denote the types of the source and target nodes, $n_{i}$ and $n_{j}$ denote their names, and $l_{i}$ denotes the edge type. ProvSyn is to learn a distribution over such provenance graphs from real system audit data, enabling the generation of synthetic graphs that preserve both the structural and semantic properties of the original dataset. The generated graphs are also heterogeneous, with diverse node and edge types, and each node is assigned a detailed name.
The objectives of ProvSyn are threefold: (1) to synthesize high-fidelity provenance graphs. This is our primary goal, as high-quality provenance data is essential for training reliable detection models; (2) to mitigate data imbalance in existing provenance graph datasets. ProvSyn is designed to learn the underlying distribution of real-world data and generate samples that represent underrepresented scenarios, thereby improving data diversity; and (3) to enhance the performance of downstream intrusion detection systems. By augmenting existing datasets with ProvSyn-generated graphs, the coverage of system behaviors can be expanded, leading to improved generalization and robustness of IDS models.
Figure 1: ProvSyn Architecture. First, a heterogeneous graph generation model is employed to generate the initial structure of a provenance graph. Subsequently, the topology of the graph is refined according to predefined rules. Finally, a large language model is used to synthesize the textual attributes of the nodes within the provenance graph.
4.2 Real-World Provenance Graph Construction
We begin by parsing each log entry to extract five key elements: source node type, source node name, destination node type, destination node name, and edge type. These elements are used to construct a directed provenance graph, where each node is assigned two attributes (type and name), and each edge is labeled with a type attribute. To avoid duplication, we maintain a hash table indexed by the MD5 hash of node names.
For log formats with simple key-value structures, we extract these elements directly using predefined keys. For formats with more complex attribute representations of entities and interactions, we apply regular expressions to extract the target elements. Extracting node names is particularly challenging; to improve accuracy, we define type-specific regular expressions tailored to each node type, ensuring that a valid name is extracted for every node.
4.3 Graph Structure Generation
Existing graph generation models primarily focus on structural connectivity, often overlooking the modeling of node and edge labels [12, 71]. Many approaches also depend on domain-specific heuristics (e.g., those tailored to molecular graphs), limiting their generalizability across graph types [56]. In contrast, provenance graphs represent a new class of heterogeneous graphs, where nodes correspond to system entities such as processes, files, and network components, and edges represent relations such as read, write, clone, and send. Generating such graphs requires a domain-agnostic model that captures both structural and label information.
To this end, we adopt GraphGen [17], a domain-independent graph generation framework capable of jointly modeling node and edge labels along with structural connectivity. GraphGen converts graphs into sequences using minimum DFS codes, a canonical representation that preserves both topology and semantics. An Long Short-Term Memory (LSTM)-based model is trained to learn the joint distribution over these sequences, enabling the generation of labeled heterogeneous graphs in a domain-agnostic fashion.
However, GraphGen is typically trained on small graphs with around 100 nodes, whereas provenance graphs often exceed 10,000 nodes, making it infeasible to feed the entire graph into the model. To make the training process scalable to large graphs, we adopt a random walk sampling strategy to extract subgraphs, and then the subgraphs can be used to construct a GraphGen-compatible training dataset. During sampling, we perform a restart-based random walk. At each step, there is a probability of returning to the starting node; otherwise, the next node is uniformly selected from the current node’s neighbors. Visited nodes and their connecting edges are dynamically added to the subgraph. The walk terminates when either a predefined number of iterations is reached or the subgraph exceeds the maximum allowed number of nodes or edges. The number of walks per starting node is determined by its degree, calculated as $\texttt{factor}\times\sqrt{\texttt{degree}}$ , ensuring that high-degree nodes generate more subgraphs. After sampling, each subgraph is normalized by remapping node IDs and removing self-loops. Only subgraphs that meet the node and edge limits and remain connected are retained. The reason why we adopt restart-based random walk for subgraph sampling is its ability to balance local exploration with broader graph coverage. Unlike simple random walk, which tends to get trapped in densely connected areas, and Metropolis-Hastings random walk, which often mixes slowly and introduces sampling bias in heterogeneous graphs, restart-based random walk periodically returns to the starting node. This mechanism helps it escape local regions and sample subgraphs that better reflect the overall graph structure. The complete procedure is presented in Algorithm 1.
The training process of GraphGen involves converting graphs into minimum DFS codes, followed by learning the distribution of these sequences using a LSTM network. Given a graph dataset $\mathcal{G}=\{G_{1},G_{2},\dots,G_{m}\}$, each graph $G_{i}$ is first transformed into its minimum DFS code sequence $S_{i}=F(G_{i})=[s_{1},s_{2},\dots,s_{|E_{i}|}]$, where $|E_{i}|$ denotes the number of edges in graph $G_{i}$, and each $s_{t}$ is an edge tuple representing the $t$-th edge in the DFS traversal.
Each edge tuple $s_{t}$ is defined as:
$s_{t}=(t_{u},t_{v},L_{u},L_{e},L_{v})$
where $t_{u}$ and $t_{v}$ are the timestamps (DFS indices) of the source and target nodes, respectively; $L_{u}$ and $L_{v}$ are the labels of the source and target nodes, respectively; and $L_{e}$ is the label of the edge between the two nodes.
At each time step $t$, the LSTM updates its hidden state $h_{t}$ based on the embedding of the previous edge tuple $s_{t-1}$, and then predicts the distribution over the next edge tuple:
$h_{t}=f_{\text{trans}}(h_{t-1},f_{\text{emb}}(s_{t-1}))$
$s_{t}\sim p(s_{t}\mid h_{t})$
Here, $f_{\text{emb}}$ is the embedding function that maps edge tuples to a vector space, and $f_{\text{trans}}$ is the transition function implemented by the LSTM.
The training objective is to minimize the binary cross-entropy (BCE) loss between the predicted and ground-truth edge tuples across the entire dataset:
$\mathcal{L}=\sum_{i=1}^{m}\sum_{t=1}^{|E_{i}|}\text{BCE}(p(s_{t}\mid s_{<t}),s_{t})$
During inference, GraphGen generates new graphs using the trained LSTM model. Starting from an initial hidden state $h_{0}$ and a start-of-sequence symbol SOS, the model iteratively generates a minimum DFS code sequence $S=[s_{1},s_{2},\dots,s_{|E|}]$. The process of updating the hidden state and sampling the next edge tuple $s_{t}$ is the same as in the training phase. The generation continues until an end-of-sequence symbol EOS is produced. The resulting DFS code sequence $S$ is then converted back into a graph $G$ by reconstructing the edges based on the timestamps and labels in the sequence, yielding the generated graph structure and labels. This approach allows the model to generate graphs that reflect the structural and semantic patterns present in the training data.
4.4 Rule-based Post-processing
After the first stage, we obtain an undirected heterogeneous graph, where both nodes and edges have distinct types. However, this graph is not directly suitable as a provenance graph structure for two main reasons. First, provenance graphs are directed by nature, whereas the generated graph is undirected. Second, despite being trained on a large dataset, the heterogeneous graph generation model occasionally produces invalid connections—such as a process-read-process relation—which do not occur in real-world scenarios.
To ensure the fidelity of generated provenance graphs, we implement a rule-based post-processing pipeline. First, we convert the undirected graph into a directed graph by replacing each undirected edge $(u,v)$ with two directed edges, $(u,v)$ and $(v,u)$. We then remove edges violating domain-specific constraints, such as connections between incompatible entity types or semantically invalid interactions. Finally, we remove all isolated nodes resulting from edge removal, as such nodes do not appear in real-world provenance graphs.
The domain-specific constraints applied during post-processing vary by dataset. For example, for the Nodlink dataset, the constraints are as follows: (1) all operations must originate from process nodes; (2) file operations must terminate at file nodes; (3) process operations must terminate at process nodes; and (4) network operations must terminate at either network or file nodes. For the Cadets, Theia, and Trace datasets, the constraints are: (1) non-read operations must originate from process nodes; and (2) read operations (e.g., read, receive, load) must terminate at process nodes.
[…]
Output : Valid subgraph $G_{\text{sampled}}$ or $\emptyset$ if invalid
1
[…]
4
5for $t\leftarrow 1$ to $T$ do
6       $r\sim\text{Uniform}[0,1]$
7       if $r<\alpha$ then
[…]
              // Restart mechanism
8
9      else
[…]
11             if $v\notin V(G_{\text{sampled}})$ then
12                   Add $v$ to $G_{\text{sampled}}$ (copy attributes from $G$)
13
[…]
16
[…]
             break
              // Node limit reached
18
[…]
             break
              // Edge limit reached
20
21Remove self-loops from $G_{\text{sampled}}$
22 Relabel nodes to consecutive integers
23 $n\leftarrow|V(G_{\text{sampled}})|$
24 $e\leftarrow|E(G_{\text{sampled}})|$
25
[…]
       return $\emptyset$
        // Subgraph is invalid
27
return $G_{\text{sampled}}$
  // Return valid subgraph
28
Algorithm 1 Restart-Based Random Walk Subgraph Sampling
4.5 Textual Attribute Generation
After the first two stages, we obtain a directed heterogeneous graph in which nodes of different types are connected by edges of different types, and all connections conform to predefined rules. However, this graph is not yet sufficient to serve as a complete provenance graph. In real-world scenarios, nodes typically carry detailed names, such as command lines or file paths. These names are essential, as many provenance-based intrusion detection methods rely on features extracted from node names using natural language processing techniques [53, 37]. Therefore, assigning meaningful node names is crucial for supporting accurate detection.
In the third stage, we leverage the natural language generation capabilities of large language models to assign names to nodes in provenance graphs. This process involves two main challenges. First, prior studies suggest that LLMs are not inherently effective at understanding graph structures  [33]. Thus, a key issue is how to transform the structure of a provenance graph into a format that is more interpretable by LLMs. Second, LLMs lack prior knowledge of typical node names in provenance graphs, and prompting them directly often results in overly simplistic outputs with limited diversity.
To address the first challenge, prior work has shown that LLMs are less effective at processing graph-structured data but perform well on sequence-based inputs [69]. Motivated by this, we propose a method that converts directed heterogeneous graphs into sequences of node and edge types using depth-first search. These sequences are then input to the LLM to generate names for the corresponding nodes. Specifically, to extract complete and non-overlapping DFS paths, we use all nodes with zero in-degree as starting points for DFS traversal. However, due to the presence of self-loops in the provenance graph, some paths may not be captured using this strategy alone. To address this, we additionally include any node involved in a self-loop as a starting point. In each extracted sequence, a node is represented as a pair $(\tau_{i},n_{i})$, where $\tau_{i}\in\mathcal{T}$ denotes the node type and $n_{i}\in\mathcal{N}$ the node name. Each edge $e_{j}$ represents the interaction type between adjacent nodes. A DFS-derived sequence is defined as:
$S=[(\tau_{1},n_{1})\xrightarrow{e_{1}}(\tau_{2},n_{2})\xrightarrow{e_{2}}\cdots\xrightarrow{e_{k-1}}(\tau_{k},n_{k})]$
For nodes without names, we use [null] as a placeholder. During the name generation process, we do not input all sequences to the LLM in parallel. This is because nodes may appear in multiple sequences, and parallel processing could lead to inconsistencies in assigned names. Instead, we sequentially feed each DFS-derived sequence to the LLM. Once a node name is generated, it is filled into the graph, ensuring that subsequent traversals can retrieve the updated name.
To address the second challenge, we construct a training dataset to provide the LLM with prior knowledge about entity names in provenance data. Each training instance is a question–answer pair, where the question (Q) consists of a DFS sequence of node and edge types, and the answer (A) is the same sequence with node names annotated. All training instances are derived from real-world provenance graphs.
To generate realistic Q&A pairs, we adopt a masking strategy that preserves structural context while requiring the model to infer node names. During the name generation process, two types of sequence are inputted to the LLM. The first type consists of sequences in which all nodes are unnamed, serving as a cold start and requiring the LLM to infer node names based solely on the order of node and edge types. The second type includes sequences where some nodes are already named, requiring the LLM to infer the remaining names using both structural order and existing node names as context. Accordingly, we define two corresponding Q&A pair formats:
Fully Masked Q&A Pairs In this setting, all node names are masked with [null] while retaining node types and edge types, creating a template for cold-start name generation. This design encourages the model to learn a mapping $f:(\tau_{i},\{\tau_{j},e_{l}\}_{j=1}^{k})\rightarrow n_{i}$, while preserving structural consistency through fixed sequences of node and edge types.
[…]
$A_{\text{full}}$ $=[(\tau_{1},n_{1})\xrightarrow{e_{1}}(\tau_{2},n_{2})\xrightarrow{e_{2}}\cdots\xrightarrow{e_{k-1}}(\tau_{k},n_{k})].$
Partially Masked Q&A Pairs To simulate realistic generation scenarios where some node names are known from prior context, we randomly mask a subset $\mathcal{M}\subset\{1,\ldots,k\}$ of node names with a masking rate $\rho$:
$Q_{\text{part}}=[(\tau_{1},\tilde{n}_{1})\xrightarrow{e_{1}}\cdots\xrightarrow{e_{k-1}}(\tau_{k},\tilde{n}_{k})],$
where
$\tilde{n}_{i}=\begin{cases}\text{[null]}&\text{if }i\in\mathcal{M},\\ n_{i}&\text{otherwise}.\end{cases}$
The corresponding answer sequence $A_{\text{part}}=A_{\text{full}}$ provides full supervision of node names. This adaptive masking strategy allows the model to learn to predict masked node names based on both the observed names and the structural context:
$P(n_{\mathcal{M}}\mid n_{\neg\mathcal{M}},\{\tau_{j}\}_{j=1}^{k},\{e_{l}\}_{l=1}^{k-1}).$
By constructing these two types of QA pairs, we provide the LLM with a comprehensive training dataset that helps it learn the correspondence between sequences and node names, thereby improving its ability to generate accurate node names in provenance graphs. This approach leverages the strengths of LLMs in sequence-based tasks while addressing the challenges of graph structure understanding, leading to more effective node name generation in provenance graphs.
4.6 Semantic Correctness Validation
Evaluating semantic correctness is essential for assessing the quality of synthetic provenance graphs, as it indicates whether the generated graphs accurately reflect the underlying logic and intent of real-world system behaviors. This evaluation examines whether the relationships among entities and the sequences of interactions conform to meaningful and valid operational patterns. Despite its importance, evaluating semantic correctness remains difficult due to the lack of standardized metrics and the inherent complexity in defining ground-truth semantics for diverse and dynamic system activities.
Model-based Evaluation Mechanism. To evaluate semantic correctness, we formally define a semantic unit in the graph as a triplet: the starting node name $s$, edge type $e$, and ending node name $t$. This unit is represented as $(s,e,t)$, where $s,t\in\mathcal{N}$ (the set of all node names) and $e\in\mathcal{E}$ (the set of all edge types). Concatenating these elements forms a natural language sentence, denoted $\text{Sentence}(s,e,t)$. This sentence captures the semantics of the interaction between entities, expressing that the source node $s$ performs an action of type $e$ on the target node $t$.
To assess the semantic correctness of $\text{Sentence}(s,e,t)$, we employ a BERT(Bidirectional Encoder Representations from Transformers) [13]-based encoder to obtain its embedding. This embedding captures the semantic representation of the interaction described by the triplet and serves as the basis for subsequent classification-based evaluation. Formally, let $\text{BERT}:\mathcal{S}\to\mathbb{R}^{d}$ be the encoding function that maps a sentence $\text{Sentence}(s,e,t)\in\mathcal{S}$ (the set of all possible semantic triplets in sentence form) to a dense vector $\mathbf{v}\in\mathbb{R}^{d}$. The embedding is computed as:
$\mathbf{v}=\text{BERT}(\text{Sentence}(s,e,t))$
However, embedding of $\text{Sentence}(s,e,t)$ captures only pairwise semantic relationships, resulting in a localized evaluation. To achieve graph-level assessment, we employ a multi-layer Graph Attention Network (GAT) [60]. This architecture aggregates neighborhood information through stacked self-attention layers, where each layer updates node representations by adaptively weighting features from adjacent nodes. We employ BERT to represent the text attributes of individual nodes as the initial node embeddings. Formally, for each node $v$ with text attribute $T_{v}$, its initial embedding $\mathbf{h}_{v}^{(0)}$ is derived via BERT:
$\mathbf{h}_{v}^{(0)}=\text{BERT}(T_{v})$
Let $\text{GAT}:\mathcal{V}\to\mathbb{R}^{d^{\prime}}$ denote the encoder that maps a node to its global representation in $\mathbb{R}^{d^{\prime}}$ using these initial embeddings. For a semantic unit $(s,e,t)$, we obtain the global embeddings of both the starting node $s$ and the ending node $t$:
$\mathbf{u}_{s}=\text{GAT}\left(s;\left\{\mathbf{h}_{v}^{(0)}\right\}_{v\in\mathcal{V}}\right),\quad\mathbf{u}_{t}=\text{GAT}\left(t;\left\{\mathbf{h}_{v}^{(0)}\right\}_{v\in\mathcal{V}}\right)$
We then concatenate the global embeddings $\mathbf{u}_{s}$ and $\mathbf{u}_{t}$ with the local embedding $\mathbf{v}$ of the semantic unit to form a comprehensive representation:
$\mathbf{z}=[\mathbf{v};\mathbf{u}_{s};\mathbf{u}_{t}]$
This comprehensive representation $\mathbf{z}$ captures both the local and global context of the semantic unit. Finally, we feed $\mathbf{z}$ into a multilayer perceptron (MLP) discriminator to determine the correctness of the semantic unit. The MLP discriminator is defined as $\text{MLP}:\mathbf{R}^{d+2d^{\prime}}\to[0,1]$. The correctness score of the semantic unit is computed as:
Correctness $\text{Correctness Score}=\text{MLP}(\mathbf{z})$
The correctness score is a value between 0 and 1, where a higher score indicates higher semantic correctness.
Contrastive Learning-based Training Framework. To train the GAT and MLP discriminator, we employ a contrastive learning approach. We use real provenance graphs as the training set, extracting all semantic units from them and labeling them as positive samples. Additionally, to provide negative samples for contrastive learning, we construct them from the positive samples using three distinct strategies to help the classifier model learn to distinguish between semantically correct and incorrect instances. The strategies are:
Subject-Object Inversion (SOI): We swap the names of the start node and the end node. This trains the model to recognize cases where the subject and object are inverted in semantics. Formally, for a semantic unit $(s,e,t)$, we generate a negative sample by inverting the subject and object:
$(s,e,t)\rightarrow(t,e,s)$
Predicate Replacement (PR): We replace the edge type with another type of edge. This trains the model to identify incorrect predicates in semantics. However, in this method, replacing the edge type with another type may still result in semantically correct cases. To guarantee the constructed semantic units are semantically incorrect, we manually classify edge types into disjoint categories. Let $\mathcal{E}=\bigcup_{i=1}^{k}\mathcal{E}_{i}$ be the partition of edge types into $k$ categories. When replacing an edge type $e\in\mathcal{E}_{i}$, we ensure the new edge type $e^{\prime}$ is from a different category, i.e., $e^{\prime}\in\mathcal{E}_{j}$ where $i\neq j$. For a semantic unit $(s,e,t)$, we generate a negative sample by replacing the predicate:
$(s,e,t)\rightarrow(s,e^{\prime},t)\quad\text{where }e^{\prime}\in\mathcal{E}_{j}\text{ and }j\neq i$
Entity Substitution (ES): We replace the name of either the start node or the end node with the name of another node. This trains the model to recognize cases where irrelevant entities appear in semantics. It is important to note that in this construction method, replacing one element of a semantic triple with another element may still result in correct semantics. This is because the replaced element and the new element may be similar. Therefore, to construct negative samples, we first cluster the node names using K-means into disjoint clusters. Let $\mathcal{V}=\bigcup_{i=1}^{m}\mathcal{V}_{i}$ be the partition of node names into $m$ clusters. When replacing a node name $s\in\mathcal{V}_{i}$ or $t\in\mathcal{V}_{i}$, we select a node $s^{\prime}\in\mathcal{V}_{j}$ or $t^{\prime}\in\mathcal{V}_{j}$ where $i\neq j$. For a semantic unit $(s,e,t)$, we generate negative samples by substituting either the subject or the object:
$(s,e,t)$ $\rightarrow(s^{\prime},e,t)\quad\text{where }s^{\prime}\in\mathcal{V}_{j}\text{ and }j\neq i$
$\text{or}\quad(s,e,t)$ $\rightarrow(s,e,t^{\prime})\quad\text{where }t^{\prime}\in\mathcal{V}_{j}\text{ and }j\neq i$
To maintain a balanced training set, we generate one negative sample for each positive sample in the real graph. Specifically, we randomly select one of the three strategies (SOI, PR, ES) to construct the negative sample. This ensures that positive and negative samples each account for 50% of the training data. For the MLP discriminator, which determines the semantic correctness of each triple, we adopt a binary cross-entropy loss. Given a batch of semantic units $\{(s_{i},e_{i},t_{i})\}_{i=1}^{N}$ with corresponding labels $y_{i}\in\{0,1\}$, the loss function is defined as:
$\mathcal{L}_{\text{MLP}}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log f_{\text{MLP}}(\mathbf{z}_{i})+(1-y_{i})\log\left(1-f_{\text{MLP}}(\mathbf{z}_{i})\right)\right]$
5 Evaluation
To thoroughly evaluate ProvSyn, we address the following research questions. Experiments on provenance graph structure generation were conducted on a machine equipped with Intel Xeon Silver CPU, 128 GB RAM, NVIDIA RTX 3090 GPUs, and Ubuntu 22.04.4 LTS. Experiments on text attribute synthesizing were performed on a machine with Intel Xeon Platinum CPU, 512 GB RAM, NVIDIA RTX A6000 GPUs, and Ubuntu 22.04.3 LTS. The research questions are:
RQ1: What is the fidelity of synthetic provenance graphs in terms of structural, textual, temporal, and embedding-based characteristics?
RQ2: What is the semantic correctness of synthetic provenance graphs?
RQ3: Can synthetic provenance graphs mitigate class imbalance in the dataset?
RQ4: When synthetic provenance graphs are used for training set augmentation in downstream intrusion detection tasks, what is the performance of the detection task?
RQ5: What are the time and resource costs of generating synthetic provenance graphs?
Further aspects of our study, including hyperparameter analysis, ablation study are discussed extensively in the Appendix C D.
5.1 Experimental Setup
Datasets. We evaluate the effectiveness of ProvSyn on two public datasets: the Nodlink Ubuntu dataset [37] and the DARPA Engagement 3 datasets [2]. The Nodlink dataset [37] we use is a simulated provenance dataset collected on Ubuntu 20.04 using Sysdig. It provides both benign logs and attack logs containing a complete multi-stage attack chain. The detailed statistics of the Nodlink provenance graphs used for training and evaluation are shown in Table I. The DARPA E3 datasets [2] are collected from an enterprise network during adversarial engagements as part of the DARPA Transparent Computing program. APT attacks are launched by a red team, while blue teams attempt to detect them through host auditing and causality analysis. TRACE, THEIA and CADETS sub-datasets are included in our evaluation. For E3 dataset, we utilize labels from ThreaTrace. As shown in Table I, the E3 dataset, compared to Nodlink, has a larger number of nodes and edges, and also exhibits greater diversity in node and edge types.
TABLE I: Comparison of dataset statistics.
Dataset Nodes Edges Node Types Edge Types
Nodlink 16,818 30,531 3 13
Cadets 89,207 276,883 6 20
Theia 77,071 202,517 3 14
Trace 74,241 113,808 7 20
Baselines. In our approach, we first use a heterogeneous graph generation model to construct the structure of the provenance graph and then employ an LLM to generate the textual attributes of the graph. In contrast, an alternative method directly calls an LLM to synthesize provenance logs and then converts these logs into a graph. This method does not require an additional heterogeneous graph generation model. We systematically evaluate and compare the provenance data synthesized by these strong LLMs with the data synthesized by ProvSyn.
To ensure a comprehensive comparison, we selected baseline models encompassing both large-parameter closed-source models and smaller-parameter open-source models. For large-parameter LLMs, we chose GPT-4o [4], Claude-3.7-Sonnet [1], and Doubao-pro-32k [3]. For smaller-parameter LLMs, we chose Qwen2.5-7B-Instruct [66] and Llama3.1-8B-Instruct [18].
To generate provenance logs using LLMs listed above, we designed structured prompts that included task descriptions, illustrative examples, and clearly defined output formats. Each dataset was paired with a customized prompt, primarily differing in the examples provided. These examples, formatted as JSON or TSV entries, contained five elements: source node type, source node name, destination node type, destination node name, and edge type. They were selected from real audit logs to ensure both representativeness and coverage of various node and edge types. For each dataset, we collected 100 responses from each LLM. After obtaining the responses, we parsed them line by line and filtered out entries that did not conform to the expected JSON or TSV format. We also discarded entries containing node or edge types outside the predefined type sets. The remaining entries were then used to construct provenance graphs for comparison with graphs generated by ProvSyn.
Configurations. For training the heterogeneous graph generation model, we sampled 20,000 subgraphs from the real provenance graphs using a random walk strategy. Among them, 16,000 subgraphs were used for training, 2,000 for validation, and 2000 for test. The model was trained for 3,000 epochs with a batch size of 32 and a learning rate of 0.003. After training, we selected the model that achieved the best performance on the validation set for use in ProvSyn. For example, on the Theia dataset, the best-performing model was obtained after 2,380 epochs and was used in the final ProvSyn pipeline.
For the text attribute synthesizing task, we used Llama3.2-3B-Instruct as the base model. We constructed a training set of 20,000 samples, consisting of 10,000 fully masked QA pairs and 10,000 partially masked QA pairs. During training, we applied the LoRA fine-tuning method, freezing most of the model parameters and updating only 0.81% of them. The model was trained for 60 steps with a learning rate of 2e-4. During inference, the temperature was set to 1.5, and the maximum number of generated tokens was set to 2048.
5.2 Fidelity Evaluation
TABLE II: Comparison of MMD metrics across different models and datasets. Lower values indicate better performance.
Dataset Model Degree Orbit Node Label Edge Label Node Label & Degree
Nodlink GPT-4o 0.75 0.94 0.11 0.75 0.89
Claude-3.7 0.56 0.85 0.07 0.92 0.77
Doubao-pro 0.58 0.82 0.18 1.50 0.59
Qwen2.5-7B 0.43 0.59 0.05 0.94 0.40
Llama3.1-8B 0.62 - 0.10 0.91 0.81
ProvSyn 0.07 0.07 0.002 0.01 0.07
Cadets GPT-4o 0.82 0.28 0.37 0.90 0.48
Claude-3.7 0.76 0.36 0.11 1.33 0.83
Doubao-pro 0.73 0.25 0.27 1.31 0.66
Qwen2.5-7B 0.76 0.39 0.67 0.77 0.33
Llama3.1-8B 0.88 0.40 0.72 0.85 0.81
ProvSyn 0.08 0.32 0.003 0.01 0.12
Theia GPT-4o 0.94 0.35 0.46 1.34 0.53
Claude-3.7 0.92 0.85 0.18 1.44 0.81
Doubao-pro 0.82 0.67 0.46 1.86 0.89
Qwen2.5-7B 0.91 0.46 0.48 1.00 0.61
Llama3.1-8B 0.68 0.87 0.53 1.01 0.66
ProvSyn 0.16 0.54 0.005 0.01 0.29
Trace GPT-4o 0.66 1.04 0.80 1.07 0.54
Claude-3.7 0.77 1.10 0.68 1.74 0.80
Doubao-pro 0.75 1.06 0.57 1.12 0.51
Qwen2.5-7B 0.24 0.23 0.44 1.06 0.47
Llama3.1-8B 0.25 0.87 1.04 1.16 0.63
ProvSyn 0.17 0.23 0.02 0.04 0.33
Structural Evaluation. To quantitatively evaluate the structural fidelity of generated graphs, we employ five Maximum Mean Discrepancy (MMD)-based metrics [19, 59] that measure distributional divergence between generated graphs and real graphs. These metrics include: (1)degree MMD for local connectivity patterns, (2)orbit MMD for higher-order structural motifs, (3)node label MMD for vertex attribute distributions, (4)edge label MMD for edge label distributions, and (5)joint node label-degree MMD for attribute-topology interactions.
To bridge the characterization of graph features with distributional divergence measurement, we leverage kernel methods to project these heterogeneous graphs into comparable representations. We construct kernel similarity matrices using two distinct Gaussian kernel formulations. The Earth Mover’s Distance-optimized Gaussian kernel is applied to categorical distribution metrics including degree MMD, node/edge label MMD, and joint node label-degree MMD, while the Euclidean-based Gaussian kernel handles continuous feature spaces in orbit MMD. The MMD² values are computed via the unbiased estimator:
$\text{MMD}^{2}=$ $\mathbb{E}_{x,x^{\prime}\sim P}[k(x,x^{\prime})]+\mathbb{E}_{y,y^{\prime}\sim Q}[k(y,y^{\prime})]$
$-2\mathbb{E}_{x\sim P,y\sim Q}[k(x,y)]$
where $P$ and $Q$ denote reference and generated graph distributions, $x,x^{\prime}$ are samples from $P$, and $y,y^{\prime}$ from $Q$.
In the experimental results shown in Table II, ProvSyn consistently outperformed the baselines across all datasets. In Nodlink dataset, ProvSyn achieved values an order of magnitude lower than the baselines on all metrics, indicating superior structural fidelity. In Cadets, Theia, and Trace datasets, ProvSyn also maintained a clear lead, particularly on label-related metrics. In contrast, baseline methods exhibit inconsistent performance patterns across the four datasets, suggesting fundamental limitations in structural generation capabilities regardless of the scale of model parameters. This finding underscores the necessity of an additional heterogeneous graph generation network, which produces structurally superior graphs compared to those generated solely by an LLM.
Texutual Evaluation. To evaluate the accuracy of generated text attributes, we use standard NLP metrics that capture different types of similarity. These include: (1) BLEU [51], which measures n-gram precision; (2) GLEU [49], which balances precision and recall; and (3) ROUGE [40], which focuses on recall by assessing how much of the reference text is covered.
For each node type, we construct a reference corpus $\mathcal{R}_{t}=\{r_{1},r_{2},...,r_{n}\}$ by aggregating all valid node names of type $t$ from the real-word provenance graph. When evaluating a generated name $g_{i}$, we compute its similarity with every reference $r_{j}\in\mathcal{R}_{t}$ through parallel metric calculations, then select the maximum similarity score as the final assessment:
$S(g_{i})=\max_{r_{j}\in\mathcal{R}_{t}}\text{sim}(g_{i},r_{j})$
The choice of maximum similarity over average similarity is driven by the following considerations. Focusing on the highest match ensures that the generated text is evaluated based on its optimal alignment with the reference corpus, thereby highlighting its capacity to produce accurate and relevant attributes. The complete procedure is presented in Algorithm 2 in Appendix B.
In our experiments, we used the Python NLTK library to compute the BLEU, GLEU, and ROUGE scores for the process and file nodes. We didn’t evaluate the text quality of network nodes because their names are purely numerical, making natural language metrics inappropriate. Additionally, our analysis shows that the network node names generated by both the baselines and ProvSyn conform to the IPv4 address format.
The results in Table III show that the node names generated by the baselines generally achieve scores below 0.1, indicating that large language models lack prior knowledge of the names of operating system entities. In contrast, ProvSyn was exposed to real entity names from audit logs during the finetuning stage, enabling a smaller 3B model to generate more accurate and realistic names. Interestingly, we observed that Llama-8B performed relatively well in the Cadets datasets. Further analysis revealed that many of the node names generated by Llama were directly copied from the examples in the prompt, even though the corresponding log entries did not make sense.
Temporal Evaluation. A provenance graph is a directed graph in which directed sequences also represent temporal sequences of events occurring in chronological order. For synthetic provenance graphs, accurately preserving the temporal structure of event sequences is essential for ensuring fidelity. To evaluate the temporal consistency of synthetic provenance graphs, we employ two metrics: Longest Common Subsequence (LCS) [7] and Dynamic Time Warping (DTW) [48].
LCS measures the length of the longest sequence of events that appears in both the real and synthetic graphs in the same order. LCS is a strict matching metric—it requires exact alignment of events and their temporal order without allowing for any deviations or shifts. DTW, on the other hand, is a more flexible metric that allows for non-linear alignments between sequences. Unlike strict matching methods, DTW allows sequences to be stretched or compressed along the time axis, thereby accommodating local variations. This flexibility makes DTW particularly suitable for scenarios where the synthetic graph captures the general pattern of the real sequence but exhibits slight shifts or inconsistencies.
In our experiments, we used depth-first search to extract sequences from the graphs. Each sequence consisted of node types and edge types, e.g. (process, clone, process, read, file). For LCS computation, we used a dynamic programming approach, and for DTW, we used the dtw library in Python. To compare synthetic graphs with the real graphs in the test set, we adopted a strategy similar to Algorithm 2 in Appendix B: each sequence from the synthetic graph was compared with all sequences from the real graph, and the best match was recorded as its score. The overall graph score was then computed by aggregating the scores of all sequences in the synthetic graph. The results in Table IV show that ProvSyn achieves higher LCS scores and lower DTW values than the baselines across all four datasets, indicating its advantage in preserving temporal consistency.
TABLE III: Comparison of textual quality across different models, types and datasets. Lower values indicate better performance.
Dataset Type Model BLEU GLEU ROUGE-L-F
Nodlink Process GPT-4o 0.05 0.11 0.23
Claude-3.7 0.08 0.15 0.27
Doubao-pro 0.02 0.03 0.09
Qwen2.5-7B 0.07 0.15 0.29
Llama3.1-8B 0.12 0.17 0.33
ProvSyn 0.89 1.00 0.99
File GPT-4o 0.09 0.11 0.29
Claude-3.7 0.10 0.11 0.29
Doubao-pro 0.06 0.11 0.25
Qwen2.5-7B 0.11 0.17 0.41
Llama3.1-8B 0.26 0.26 0.37
ProvSyn 0.31 0.32 0.44
Cadets Process GPT-4o 0.24 0.24 0.24
Claude-3.7 0.17 0.17 0.17
Doubao-pro 0.03 0.03 0.03
Qwen2.5-7B 0.03 0.03 0.04
Llama3.1-8B 1.00 1.00 1.00
ProvSyn 1.00 1.00 1.00
File GPT-4o 0.01 0.01 0.20
Claude-3.7 0.03 0.03 0.14
Doubao-pro 0.01 0.01 0.14
Qwen2.5-7B 0.02 0.02 0.16
Llama3.1-8B 0.00 0.00 0.03
ProvSyn 0.31 0.32 0.60
Theia Process GPT-4o 0.08 0.08 0.09
Claude-3.7 0.04 0.04 0.05
Doubao-pro 0.01 0.01 0.03
Qwen2.5-7B 0.19 0.19 0.20
Llama3.1-8B 0.00 0.00 0.00
ProvSyn 0.76 0.89 0.47
File GPT-4o 0.03 0.03 0.21
Claude-3.7 0.05 0.05 0.16
Doubao-pro 0.01 0.01 0.11
Qwen2.5-7B 0.00 0.00 0.29
Llama3.1-8B 0.00 0.00 0.35
ProvSyn 0.37 0.44 0.48
Trace Process GPT-4o 0.02 0.02 0.02
Claude-3.7 0.02 0.02 0.02
Doubao-pro 0.00 0.00 0.00
Qwen2.5-7B 0.02 0.02 0.02
Llama3.1-8B 0.00 0.00 0.00
ProvSyn 0.50 0.50 0.50
File GPT-4o 0.00 0.00 0.03
Claude-3.7 0.00 0.00 0.03
Doubao-pro 0.00 0.00 0.01
Qwen2.5-7B 0.00 0.00 0.02
Llama3.1-8B 0.00 0.00 0.00
ProvSyn 0.00 0.00 0.67
Graph Embedding Evaluation. To evaluate the fidelity of synthetic graphs from an embedding perspective, we employ graph representation learning techniques to project both synthetic and real graphs into a shared vector space, and measure their similarity using cosine similarity. Specifically, we adopt two representative embedding methods: DeepWalk [52] and Doc2Vec [35].
DeepWalk is a foundational algorithm in graph representation learning. It learns node embeddings by simulating truncated random walks on the graph and treating the resulting node sequences as sentences for a language modeling task. Given a graph $G$, we perform $N$ random walks of length $L$ starting from each node, generating a corpus $\mathcal{W}=\{w_{1},w_{2},\ldots,w_{N}\}$, where each $w_{i}$ is a walk sequence of node identifiers. These sequences are then used to train a Word2Vec model, producing node-level embeddings. The graph-level embedding $\mathbf{v}_{\text{deepwalk}}$ is computed as the mean of all node embeddings:
$\mathbf{v}_{\text{deepwalk}}=\frac{1}{|\mathcal{V}|}\sum_{v\in\mathcal{V}}\text{deepwalk}(v)$
Notably, DeepWalk does not consider textual attributes of nodes.
Unlike DeepWalk, Doc2Vec leverages node name information to capture the textual semantics of the graph. It treats each graph as a document, which is constructed from a collection of random walks. For a given graph $G$, we perform $N$ random walks of length $L$, resulting in a sequence set $\mathcal{W}=\{w_{1},w_{2},\ldots,w_{N}\}$, where each $w_{i}$ is a sequence of node names. These sequences are used to train a Doc2Vec model, yielding a graph-level embedding $\mathbf{v}_{\text{doc2vec}}$:
$\mathbf{v}_{\text{doc2vec}}=\text{doc2vec}(\mathcal{W})$
To assess the similarity between a synthetic graph $G_{s}$ and a real graph $G_{r}$, we compute the cosine similarity between their respective embeddings:
$\text{Similarity}=\frac{\mathbf{v}_{G_{s}}\cdot\mathbf{v}_{G_{r}}}{\|\mathbf{v}_{G_{s}}\|\|\mathbf{v}_{G_{r}}\|}$
A higher similarity score indicates that the synthetic graph exhibits greater fidelity to the real graph in the embedding space.
For DeepWalk-based similarity, we use the Word2Vec model from the Gensim library, and for Doc2Vec-based similarity, we use Gensim’s Doc2Vec model. In our experiments results shown in Table IV, ProvSyn outperforms the baselines on the Nodlink, Cadets, and Trace datasets, while performing slightly worse than GPT and Claude on Theia. The results show that ProvSyn maintains stable embedding similarity, with DeepWalk scores above 0.80 and Doc2Vec scores above 0.65 across four datasets. In comparison, the baselines exhibit greater variation across different datasets.
TABLE IV: Comparison of Temporal and Embedding metrics across different datasets and models. For LCS, higher is better; for DTW, lower is better; for DeepWalk and Doc2Vec, higher is better.
Dataset Model Temporal Metrics Embedding Metrics
LCS DTW DeepWalk Doc2Vec
Nodlink GPT-4o 2.44 1.54 0.66 0.69
Claude-3.7 3.03 1.33 0.83 0.40
Doubao-pro 2.64 0.89 0.60 0.59
Qwen2.5-7B 2.84 0.56 0.58 0.32
Llama3.1-8B 2.88 0.83 0.50 0.16
ProvSyn 3.30 0.41 0.89 0.74
Cadets GPT-4o 2.99 0.48 0.70 0.40
Claude-3.7 2.87 1.20 0.73 0.39
Doubao-pro 2.68 3.62 0.64 0.71
Qwen2.5-7B 2.13 24.30 0.79 0.25
Llama3.1-8B 2.96 0.48 0.48 0.67
ProvSyn 3.01 0.21 0.90 0.82
Theia GPT-4o 2.39 0.65 0.89 0.53
Claude-3.7 1.80 0.74 0.82 0.70
Doubao-pro 2.78 1.43 0.56 0.39
Qwen2.5-7B 1.36 8.90 0.74 0.48
Llama3.1-8B 1.04 8.94 0.39 0.66
ProvSyn 3.02 0.00 0.81 0.65
Trace GPT-4o 2.16 0.46 0.70 0.27
Claude-3.7 2.15 0.54 0.67 0.41
Doubao-pro 2.09 0.40 0.57 0.08
Qwen2.5-7B 2.10 0.47 0.75 0.09
Llama3.1-8B 2.08 4.45 0.71 -0.01
ProvSyn 2.80 0.10 0.81 0.75
5.3 Semantic Correctness Evaluation
Figure 2: Comparison of semantic accurary across different models and datasets. Higher values indicate better performance.
We use the semantic correctness validation mechanism introduced in Section 4.6 to evaluate the semantic of synthetic graphs. In the experiment, the GAT model was trained for 2500 epochs with a learning rate of 0.0001, while the MLP classifier was trained for 1500 epochs with the same learning rate. To validate the effectiveness and interpretability of our semantic accuracy evaluation, we tested it on real-world provenance graphs from the same datasets but different from the training graphs. The results show that the semantic accuracy of these real graphs exceeds 90% across the Nodlink, Cadets, Theia, and Trace datasets. This supports the validity of our semantic accuracy evaluation method. The results in Figure 2 show that the provenance graphs generated by ProvSyn consistently outperform the baselines in semantic accuracy across all four datasets, showing robustness to differences in data distribution. In contrast, the baselines exhibit substantial fluctuations in semantic accuracy across datasets. Notably, in the more challenging Theia and Trace datasets, the semantic accuracy of baseline-generated graphs remained below 0.20, whereas ProvSyn achieved 0.57 and 0.82, respectively. These results demonstrate the robustness of ProvSyn in handling complex, low signal-to-noise provenance data.
5.4 Data Imbalance Evaluation
To evaluate whether the inclusion of ProvSyn-generated synthetic graphs mitigates class imbalance, we use two metrics—entropy and the Gini index—to quantify the distributional balance of node and edge labels. Given a label distribution with probabilities $p_{1},p_{2},\ldots,p_{n}$, the two metrics are defined as:
$\text{Entropy}=-\sum_{i=1}^{n}p_{i}\log(p_{i}),\quad\text{Gini}=1-\sum_{i=1}^{n}p_{i}^{2}$
Higher values of both metrics indicate a more uniform and balanced label distribution.
We conduct experiments on the Nodlink, Cadets, Theia, and Trace datasets, both in their original data and after incorporating ProvSyn-generated data. For each dataset, we add 100 synthetic graphs merged with the original graph as graph communities. The results are summarized in Table V. As shown, the inclusion of ProvSyn increases both entropy and Gini index for node and edge labels across all datasets. In the Nodlink dataset, the metrics improve significantly due to the dataset’s small scale. These results suggest that ProvSyn contributes to a more balanced label distribution by enriching underrepresented classes in provenance graph datasets.
To evaluate whether the synthesized graphs extend beyond the training data distribution and contain novel structures that cover more provenance scenarios, we use two metrics: Novelty and Uniqueness. Novelty is computed by verifying that each generated graph is neither a subgraph of any reference graph in the training set nor contains any reference graph as a subgraph, indicating whether novel structures exist in the synthesized graphs. Uniqueness measures the proportion of generated graphs that are structurally distinct from one another, reflecting the diversity of the generated graphs. As shown in Table VI, the graphs generated by ProvSyn achieve 100% novelty and 100% uniqueness across all four datasets. This indicates that the synthetic data exhibits strong diversity and effectively captures patterns that were underrepresented or missing in the original data.
TABLE V: Comparison of node and edge label entropy and Gini index across different datasets with and without synthetic data. Higher values indicate better performance.
Dataset Data Source Node Label Edge Label
Entropy Gini Index Entropy Gini Index
Nodlink Real 1.03 0.42 2.41 0.75
+ProvSyn 1.31 (+0.28) 0.55 (+0.13) 2.63 (+0.22) 0.82 (+0.07)
Cadets Real 1.11 0.44 1.84 0.52
+ProvSyn 1.31 (+0.20) 0.53 (+0.09) 1.92 (+0.08) 0.54 (+0.02)
Theia Real 1.31 0.53 1.33 0.39
+ProvSyn 1.41 (+0.10) 0.58 (+0.05) 1.36 (+0.03) 0.39 (+0.00)
Trace Real 1.52 0.57 1.57 0.45
+ProvSyn 1.62 (+0.10) 0.61 (+0.04) 1.61 (+0.04) 0.47 (+0.02)
TABLE VI: Evaluation of ProvSyn-generated graphs in terms of Novelty and Uniqueness across different datasets. Higher values indicate better performance.
Dataset Novelty (%) Uniqueness (%)
Nodlink 100.0 100.0
Cadets 100.0 100.0
Theia 100.0 100.0
Trace 100.0 100.0
5.5 Application in IDS
To demonstrate the practical utility of ProvSyn, we focus on the task of APT attack detection. We conduct experiments using two detection algorithms, Magic[31] and Nodlink[37]. Specifically, we augment the training set with synthesized provenance graphs and then observe whether the detection performance improves or declines during testing.
For the Nodlink algorithm, we enhanced its dataset by adding 50 synthetic graphs merged with the original provenance graph as graph communities. This resulted in a 15% increase in node counts. For the Magic algorithm, we enhanced the Cadets, Theia, and Trace datasets. Due to their larger scale, we added 300 synthetic graphs merged into their original provenance graphs. This led to a 10% to 15% increase in node counts.
To evaluate the effectiveness of synthetic graphs in intrusion detection tasks, we maintain identical training configurations before and after incorporating synthetic data. The detection performance is presented in Table VII. In addition to Precision and Recall, we report the number of False Positives (#FP), an crucial metric for intrusion detection system, as a high false positive rate can increase the workload of subsequent security analysis. As shown in the results, on the Nodlink dataset, the inclusion of synthetic graphs leads to no change in Precision, Recall, or the number of false positives, indicating that the synthetic data is of comparable quality to the original data. On the Cadets, Theia, and Trace datasets, Recall remains constant, and Precision improves. Specifically, false positives decrease by 506 on Cadets (from 6435 to 5929), by 95 on Theia (from 931 to 836), and by 379 on Trace (from 2390 to 2011). These results suggest that augmenting with synthetic graphs produced by ProvSyn improves model performance, and effectively reduces false positives in detection systems.
TABLE VII: IDS performance across different datasets with and without synthetic data. Higher values indicate better performance.
Dataset Data Source Precision(%) Recall(%) #FP
Nodlink Real 100.00 100.00 0
+ ProvSyn 100.00 100.00 0
Cadets Real 66.58 99.79 6435
+ ProvSyn 68.37 99.77 5929
Trace Real 96.61 99.98 2390
+ ProvSyn 97.13 99.98 2011
Theia Real 96.45 99.99 931
+ ProvSyn 96.80 99.99 836
5.6 Overhead Analysis
TABLE VIII: Time and memory requirements of ProvSyn’s core modules.
Module Phase Time Memory
Graph Generation Training 18 hours 0.84 GB
Inference 28.09 s 0.38 GB
Text Generation Training 109 s 3.44 GB
Inference 2.78 hours 28 GB
In this section, we analyze the time and memory requirements of ProvSyn (see Table VIII). The most time- and resource-intensive components of the ProvSyn framework are the heterogeneous graph generation module and the text attribute generation module.
For the heterogeneous graph generation module, we use 16,000 samples for training, with a batch size of 32 and 3,000 training epochs. The average training time is 18 hours, with a GPU memory usage of 0.84 GB. Inference for generating 1,000 graphs takes 28.09 seconds and consumes 0.38 GB of memory.
For the text attribute generation module, we adopt Llama3.2-3B-Instruct as the base model, applying the LoRA algorithm with 4-bit quantization. The FastLanguageModel library is used to accelerate training. The training consists of 60 steps, takes an average of 109 seconds, and requires a peak memory of 3.44 GB. During inference, generating text attributes for a single graph takes an average of 80 seconds. In our experiments, we deploy 8 fine-tuned LLMs in parallel to synthesize 1,000 graphs, completing the process in 2.78 hours. The total GPU memory required for the 8 models is 28 GB.
This analysis demonstrates that ProvSyn is a time- and memory-efficient framework for provenance graph synthesis.
6 Discussion and Limitations
Temporal Modeling. A provenance graph is not only a directed heterogeneous graph with textual attributes, but also a temporal graph in which each event is associated with a timestamp indicating its occurrence time in the system. These temporal information has recently been leveraged by downstream intrusion detection systems [53] to better capture the patterns of benign and anomalous behaviors. However, in ProvSyn, we do not synthesize explicit timestamps for events. Instead, we approximate temporal order by relying on the sequence of directed edges. Future work may explore modeling temporal dynamics using temporal graph generation networks [75], or designing strategies that leverage LLMs to generate realistic timestamps, thereby enabling more comprehensive and temporally coherent provenance graph synthesis.
Graph Scale. In ProvSyn, the generated provenance graphs are smaller in scale compared to real-world provenance graphs. To enable data augmentation, we embed multiple synthetic graphs as distinct communities within large real-world provenance graphs. The results show that this approach effectively mitigates data imbalance and reduces the false positive rate of detection models trained on the augmented datasets. Future work can focus on scaling up synthetic provenance graphs by enhancing graph generation models for large-scale synthesis and exploring graph merging algorithms that combine smaller graphs into larger ones while preserving structural fidelity.
Efficiency of Text Attribute Synthesis. During textual attribute synthesis, overlaps among nodes in different sequences generated by DFS can lead to naming conflicts when these sequences are processed in parallel to generate node names. To address this issue, we serialize the sequences and input them into the LLM sequentially. The generated node names from earlier sequences are then used as context for subsequent sequences. This approach is effective for small-scale graphs. However, the synthesis efficiency decreases for large-scale graphs. Future work may explore strategies to partition large graphs into batches of non-overlapping sequences, enabling parallel processing while avoiding node naming conflicts.
Application Scenarios. To date, the graphs synthesized by ProvSyn have demonstrated utility within the domain of provenance graph datasets and for detection algorithms targeting APT attacks. However, ProvSyn is a general framework capable of synthesizing high-fidelity heterogeneous graphs with textual attributes. This gap between ProvSyn’s inherent capabilities and current applications presents an opportunity for future work: expanding the exploration of ProvSyn’s capabilities to encompass a wider range of application scenarios and downstream algorithms.
7 Conclusion
This paper proposes ProvSyn, a novel provenance graph generation framework consisting of three stages: heterogeneous graph structure generation, rule-based topological refinement, and LLM-based synthesis of textual attributes. To evaluate the fidelity of the generated graphs, we assess structural, textual, temporal, embedding, and semantic aspects. Compared with provenance data directly synthesized by LLMs, ProvSyn consistently produces graphs with higher fidelity. When integrated into existing datasets, ProvSyn-generated graphs help mitigate class imbalance and lead to reduced false positive rates in detection models trained on the augmented data.
Appendix A Entity and Event Distribution in Our Datasets
We present the distributions of entity type and event type across four commonly-used public provenance datasets in Figure 3. The distributions exhibit a typical long-tail pattern and reveal significant data imbalance.
Appendix B Textual Quality Evaluation
In Algorithm 2, we compute the maximum similarity score between the text attribute of each node and the reference corpus. This metric serves to evaluate the model’s maximum potential in generating accurate and relevant attributes.
Input : Generated graph nodes $V_{\text{gen}}=\{g_{1},g_{2},...,g_{m}\}$ with predicted names and types;
Real graph nodes $V_{\text{real}}=\{r_{1},r_{2},...,r_{n}\}$ with ground-truth names and types;
Node type set $\mathcal{T}=\{t_{1},t_{2},...,t_{k}\}$
Output : Final textual quality score $S_{\text{final}}$
1
2foreach $t\in\mathcal{T}$ do
       $\mathcal{R}_{t}\leftarrow\{r_{j}\mid r_{j}\in V_{\text{real}},\text{type}(r_{j})=t\}$
        // Construct reference corpus
3
[…]
6
7foreach $g_{i}\in V_{\text{gen}}$ do
8       $t\leftarrow\text{type}(g_{i})$
[…]
10       foreach $r_{j}\in\mathcal{R}_{t}$ do
[…]
14
[…]
17
[…]
  // Aggregate final score
18
19 return $S_{\text{final}}$
20
Algorithm 2 Textual Quality Evaluation
Appendix C Hyperparameters Setting
Figure 3: Entity Type and Event Type Distribution in Provenance Dataset including Nodlink, Cadets, Theia and Trace.
Figure 4: Hidden size setting in Nodlink dataset.
Figure 5: Inference temperature setting in Nodlink dataset.
Figure 6: Ablation Study.
Hidden size is of LSTM for heterogeneous graph generation. Hidden size is a key parameter for LSTM, as it determines the capacity of the model to capture temporal and structural dependencies in sequential node and edge generation. We conducted experiments on the Nodlink dataset and evaluated structural fidelity using five MMD-based metrics in the validation set. The results in Figure 4 show that a hidden size of 64 yields significantly lower fidelity compared to the other settings. For hidden sizes of 128, 256, and 512, the MMD values remain at a comparable level, with 256 performing slightly better than 128 and 512. Considering both performance and computational cost, we select 256 as the hidden size for ProvSyn.
Inference temperature is a key parameter in LLM-based inference, controlling the randomness of generated text. Higher temperatures yield more diverse outputs, while lower temperatures produce more deterministic results. We conducted experiments on the Nodlink dataset using temperature values of 0.1, 0.5, 1.0, and 1.5, with all other inference parameters held constant. Evaluation focused on the quality of text generated for process and file node attributes, as well as semantic correctness. As shown in Figure 5, the best performance is achieved when the temperature is set to 1.5. In particular, at lower temperatures, the generated node names tended to be repetitive and lacked diversity. Based on these findings, we selected 1.5 as the final inference temperature for ProvSyn, consistent with the recommendation of the FastLanguageModel library.
Appendix D Ablation Study
Alternative models to GraphGen. In the heterogeneous graph generation module, we adopt the design of GraphGen, which encodes graphs as DFS code sequences and employs an LSTM for training and inference. To validate the effectiveness of this design, we compare it with alternative models, including the widely used GraphRNN [72]. In our experiments, we evaluate the graphs generated by GraphRNN and GraphGen using five MMD metrics. As shown in Figure 6(a), the graphs generated by GraphGen consistently achieve significantly lower MMD scores across all five metrics, indicating substantially better fidelity compared to those generated by GraphRNN. These results demonstrate the effectiveness of using GraphGen for heterogeneous graph generation in ProvSyn.
Masking strategies. For text attribute generation, we employ two masking strategies to construct the training set, corresponding to two different input sequences during inference. To evaluate the effectiveness of the masking design, we perform ablation experiments by disabling one of the strategies and using only the other. Model performance is then evaluated using text quality metrics and semantic correctness. The experiments are conducted on the Trace dataset, with the number of training samples fixed at 20,000 for both the single-strategy and dual-strategy settings. In the results shown in Figure 6(b), we observe that using only the Full Masking strategy yields better performance than using only the Part Masking strategy. This may be because the Part Masking strategy provides partial node names within the sequence, offering more cues for the LLM to leverage. As a result, a model trained solely with Part Masking performs poorly when tested on Full Masking sequences, which contain fewer clues and less information. In contrast, using both strategies together achieves better performance than either alone, as it exposes the model to a broader range of sequence types that may appear during inference. These results demonstrate the effectiveness of our proposed approach that combines both masking strategies.
LLM parameter size. In our experiments on text attribute generation, we employ the Llama-3.2-3B-Instruct model as the base model for training. We also compare its performance against models of different sizes, including Llama-3.2-1B-Instruct and Llama-3.1-8B-Instruct. All models are trained on the same training datasets and evaluated using both text quality metrics and the semantic correctness metric. The results, presented in Figure 6(c), demonstrate that the 3B model achieves the highest scores on the BLEU and GLEU metrics. In contrast, the 8B model shows marginally better performance on the ROUGE and semantic correctness metrics. Given that the advantage of the 8B model is not substantial and considering computational resource constraints, we select the 3B model as the base for synthesizing text attributes. In practice, ProvSyn introduces a general training strategy to leverage LLMs for synthesizing text attributes, where the choice of the base model involves balancing the complexity of the dataset, the application requirements and the limitations of computational resources.

### v2

@@
-rtboxstyle linecolor=black, backgroundcolor=lightlight-gray, leftmargin=0cm, rightmargin=0cm, skipabove=4pt, skipbelow=4pt, linewidth=0.4pt, innertopmargin=3pt, innerbottommargin=3pt, innerrightmargin=6pt, innerleftmargin=6pt, roundcorner=2
+No Data? No Problem: Synthesizing Security Graphs for Better Intrusion Detection
 Abstract
@@
 While recent systems combine graph neural networks (GNNs) with natural language processing (NLP) to capture structural and semantic features, their effectiveness is limited by class imbalance in real-world data.
-To address this, we introduce ProvSyn, an automated framework that synthesizes provenance graphs through a three-phase pipeline: (1) heterogeneous graph structure synthesis with structural-semantic modeling, (2) rule-based topological refinement, and (3) context-aware textual attribute synthesis using large language models (LLMs).
-ProvSyn includes a comprehensive evaluation framework that integrates structural, textual, temporal, and embedding-based metrics, along with a semantic validation mechanism to assess the correctness of generated attack patterns and system behaviors.
-To demonstrate practical utility, we use the synthetic graphs to augment training datasets for downstream APT detection models.
-Experimental results show that ProvSyn produces high-fidelity graphs and improves detection performance through effective data augmentation.
+To address this, we introduce ProvSyn, a novel hybrid provenance graph synthesis framework, which comprises three components: (1) graph structure synthesis via heterogeneous graph generation models, (2) textual attribute synthesis via fine-tuned Large Language Models (LLMs), and (3) five-dimensional fidelity evaluation.
+Experiments on six benchmark datasets demonstrate that ProvSyn consistently produces higher-fidelity graphs across the five evaluation dimensions compared to four strong baselines.
+To further demonstrate the practical utility of ProvSyn, we utilize the synthesized graphs to augment training datasets for downstream APT detection models.
+The results show that ProvSyn effectively mitigates data imbalance, improving normalized entropy by up to 35%, and enhances the generalizability of downstream detection models, achieving an accuracy improvement of up to 38%.
 1 Introduction
-Data provenance [29], which involves recording the history of data modifications, access, and usage, serves as a critical tool in cybersecurity systems for tracing attacks [61].
-However, as system scales expand, the volume of provenance data increases dramatically, posing challenges for data management and analysis.
-To address this, researchers have begun to represent the provenance of the data using graph structures and applying graph machine learning algorithms for the detection and tracking of intrusions [21, 62, 53].
-This approach effectively captures the complexity of system activities and enables the detection of attacks that traditional methods struggle to identify, such as zero-day attacks and APT attacks.
-The quality of provenance data sets largely determines the effectiveness of provenance-based intrusion detection systems (IDS) trained on them.
-However, there are very few datasets available, and even fewer datasets that can accurately describe benign and malicious activities in a system.
-One possible solution is to produce system logs captured from real-world scenarios.
-However, one of the key disadvantages of using such datasets is privacy concerns.
-Sharing such dataset across the community could result in leakage of critical privacy information, and hence it’s difficult for organizations to release such datasets.
-Another solution is to collect benign and malicious activities in experimental environments.
-For example, the DARPA Transparent Computing (TC) [2] program contributes two widely used datasets in the cybersecurity community, DARPA TC dataset and DARPA OpTC dataset.
-However, such datasets, while giving us insights into system activities, convey limitations, one of which being the class imbalance issue [6], where less common or underrepresented program behaviors are under-sampled.
-This imbalance hampers the generalization ability of machine learning models trained on such data, limiting their effectiveness in unseen environments [77, 9, 76, 6].
-To mitigate the issue of data imbalance, a common approach is to synthesize high-fidelity provenance data that expands coverage of minority classes and patterns.
-Such data synthesis can be achieved through methods such as graph generation models or large language models (LLMs), yet each has critical limitations.
-On the one hand, most graph generation models focus on structural connectivity and lack support for generating rich textual attributes, which are essential for downstream intrusion detection tasks [12, 71, 37, 53].
-On the other hand, although LLMs are strong at producing semantically rich log entries, their outputs often fail to preserve the underlying graph structure when transformed into provenance graphs [4, 1].
-To overcome these challenges, we propose ProvSyn, a hybrid framework that integrates the structural modeling ability of graph generation models with the text generation ability of LLMs to jointly synthesize provenance graphs with both accurate topology and meaningful textual attributes.
-Specifically, ProvSyn is a three-phase synthesis framework comprising structural synthesis, topological refinement, and textual attribution.
-The initial phase constructs heterogeneous graph topology by generating nodes and edges with categorical attributes.
-Subsequent structural refinement applies domain-specific rules to eliminate unreasonable connections and isolated entities.
-In the third phase, an LLM is used to generate textual attributes of nodes.
-However, the implementation of our framework confronts three principal technical challenges:
+Data provenance [inam2023sok] records system activities for cybersecurity attack tracing [wang2020you].
+Recent deep learning-based pids (pids[han2020unicorn, wang2022threatrace, zengy2022shadewatcher, rehman2024flash, li2023nodlink] effectively capture system complexity and detect sophisticated attacks like zero-day and apt (apt) attacks.
+These systems typically rely on unsupervised or self-supervised learning to model normal system behaviors.
+However, their effectiveness is constrained by training dataset quality and diversity [11023260].
+These systems assume training data adequately represents diverse runtime configurations, workload patterns, and benign activities.
+However, this assumption is often violated in real-world scenarios where publicly available datasets exhibit significant limitations.
+Datasets that used for training and evaluating learning-based pids often suffer from data imbalance, where certain system behaviors are overrepresented while others are severely underrepresented.
+These datasets simulate cyber attacks primarily from controlled experimental environments such as the DARPA TC and OpTC datasets [darpa-E3].
+Prior studies show that benign activities in these datasets predominantly originate from automated workload generators, producing highly monotonous and predictable background behaviors [11023260].
+Consequently, models trained on such data tend to overfit to repetitive patterns and fail to generalize to diverse, unpredictable activities in production environments.
+This data imbalance hampers the generalizability of machine learning-based pids, limiting their effectiveness in diverse real-world deployments [zhou2023improving, chawla2002smote, zhao2021graphsmote, anjum2021analyzing, 11023260].
+To synthesize provenance graphs for data augmentation, our key insight is that provenance graphs encode two complementary types of information: structural information capturing relationships and interaction patterns among system entities, and semantic information describing entity attributes and identities.
+Both are critical for learning-based intrusion detection, as malicious activities are jointly determined by structural patterns (e.g., abnormal execution chains) and semantic attributes (e.g., suspicious file paths or command-line arguments) [li2023nodlink, rehman2024flash].
+However, existing synthesis approaches fail to handle this dual nature.
+Graph synthesis methods [you2018graphrnn, goyal2020graphgen, de2018molgan] focus on generating topologies and categorical labels while neglecting rich textual attributes essential for intrusion detection.
+Conversely, LLM-based text generation excels at producing semantically rich log entries but fails to preserve graph structure when converted to provenance graphs, resulting in structurally invalid or semantically inconsistent connections.
+Therefore, effective synthesis must jointly model both structural connectivity and semantic attributes to generate high-fidelity provenance graphs for robust detection systems.
+To address these limitations, we propose ProvSyn, a hybrid provenance graph synthesis framework integrating graph generation models and large language models.
+ProvSyn jointly models both structural and semantic information through a two-phase pipeline.
+In the graph structural synthesis phase, a heterogeneous graph generation model constructs topologies with accurate node and edge type distributions, capturing structural patterns from real provenance data.
+In the textual attribute synthesis phase, a fine-tuned LLM generates realistic textual attributes for each node, ensuring semantic consistency with the graph structure.
+However, realizing this framework introduces three principal technical challenges:
 Scalability to provenance graphs.
-Provenance graphs are heterogeneous graphs with more than 10,000 nodes and edges, whereas most heterogeneous graph generation models are designed for molecular graphs and are typically trained on graphs with fewer than 100 nodes [72, 17].
-Effectively scaling graph generation models designed for other domains to large-scale provenance graphs thus remains a challenge.
+Provenance graphs often contain millions of nodes and edges, whereas most existing heterogeneous graph generation models are trained on graphs with fewer than 100 nodes.
+To address this scalability challenge, we adopt rwr (rwr) sampling to extract representative subgraphs.
+Unlike simple random walks that become trapped in densely connected regions, rwr periodically restarts from the initial node, enabling broader exploration.
+We then train GraphGen [goyal2020graphgen] on sampled subgraphs to jointly model node types, edge types, and structural connectivity.
+To guarantee validity and realism, we apply domain-specific refinement rules that eliminate invalid connections (e.g., process-clone-network).
 Textual attributes synthesis.
-First, LLMs inherently struggle with graph structure comprehension [33], necessitating transformation of provenance graphs into LLM-interpretable formats.
-Second, due to lacking prior knowledge of typical provenance entity names, direct LLM prompting produces simplistic, low-diversity outputs.
+Although LLMs demonstrate strong text generation capabilities, they struggle with graph structure comprehension [jin2024large] and lack domain-specific provenance knowledge.
+Consequently, directly prompting LLMs often produces generic or unrealistic outputs.
+To address this, we first design a sft (sft) strategy adapting LLMs to the provenance domain.
+Then we convert provenance graphs into dfs (dfs) sequences, transforming graph structure into a linear format suitable for LLM processing while preserving connectivity and relationships.
 Evaluation framework gap.
-The absence of a multidimensional evaluation framework hinders comprehensive assessment of synthesized graphs, particularly in verifying the semantic correctness of attack patterns and system behavior within generated provenance structures.
-To address the first challenge, we adopt a domain-agnostic heterogeneous graph generation network [17] to jointly model node and edge labels along with structural connectivity.
-To make the training process scalable to large graphs, we adopt a restart-based random walk for subgraph sampling.
-Through a novel restart mechanism, this sampling algorithm escapes local regions and captures subgraphs that better reflect the overall graph structure.
-To address the second challenge, we convert the graph structure into depth-first search (DFS) sequences to enhance the LLM’s understanding of the graph.
-In addition, we also designed two different masking strategies, including fully masking and partially masking, to construct a training set for fine-tuning the LLM, thereby providing it with knowledge of entity names in the provenance graph.
-To address the third challenge, we design a comprehensive framework that evaluates fidelity from five perspectives: structural, textual, temporal, embedding, and semantic correctness.
-For structural fidelity, we adopt five Mean Discrepancy (MMD)-based metrics.
-For textual fidelity, we use standard NLP metrics, including BLEU, GLEU, and ROUGE.
-For temporal fidelity, we use sequence comparison metrics including Longest Common Subsequence (LCS) and Dynamic Time Warping (DTW).
-For embedding fidelity, we utilize various embedding methods, including DeepWalk and Doc2Vec.
-To accurately assess semantic correctness—the most challenging aspect—we adopt a model-based approach.
-Specifically, we sample positive and negative triples from real provenance graphs and train a classification model using contrastive learning.
-In the experiments, we used our proposed evaluation framework to evaluate the provenance data synthesized by LLMs, including GPT-4o and Claude 3.7, as well as by ProvSyn.
-The results show that ProvSyn achieves higher fidelity than these strong LLMs.
-Moreover, by generating 100% novel and 100% diverse graphs, ProvSyn effectively mitigates the data imbalance issue in provenance datasets.
-When applied to an intrusion detection system, models trained on datasets augmented with ProvSyn-generated graphs exhibited reduced false positive rates during testing.
+A critical challenge in provenance graph synthesis is the absence of comprehensive evaluation frameworks.
+Existing methods [you2018graphrnn, li2018learning, goyal2020graphgen], focusing on topological properties are inadequate for provenance graphs requiring verification of both structural validity and semantic plausibility.
+To address this gap, we propose a novel five-dimensional fidelity evaluation framework, including structural, textual, temporal, embedding, and semantic metrics.
+This framework enables rigorous assessment of synthesized provenance graphs and addresses a critical gap in the graph generation literature.
+In the experiments, we evaluate the fidelity of provenance data synthesized by ProvSyn following our proposed evaluation framework.
+We compare ProvSyn against four advanced LLMs (GPT-5.1 [openai-gpt51], Claude-Sonnet-4.5 [claude], DeepSeek-v3.2 [liu2025deepseek], and Qwen3-Max [yang2025qwen3]) on six widely used provenance datasets.
+The results show that ProvSyn achieves higher fidelity than directly using these LLMs from all five evaluation perspectives.
+Moreover, data imbalance is consistently mitigated after incorporating ProvSyn, improving normalized entropy by up to 35%.
+We further demonstrate the practical utility of ProvSyn by augmenting existing datasets with the synthesized graphs for improving four latest downstream APT detection models (NodLink [li2023nodlink], Magic [jia2024magic], Flash [rehman2024flash], Velox [bilot2025sometimes]).
+Experimental results demonstrate that detection models trained on ProvSyn-augmented datasets yield up to a 38% gain in accuracy, while concurrently bolstering discriminative reliability and diagnostic utility.
 The main contributions of our work are as follows:
-  • We propose ProvSyn, a three-phase provenance graph synthesis framework that enables the construction of high-fidelity heterogeneous graphs with rich textual attributes.
-  • We design a comprehensive multi-dimensional evaluation framework that evaluates the fidelity of synthesized provenance graphs, including: structural, textual, temporal, embedding, semantic correctness.
-  • We mitigate the data imbalance problem in provenance datasets by synthesizing novel and diverse provenance graphs, thereby enhancing the generalizability of downstream intrusion detection models.
-  • The proposed generation and evaluation frameworks can also be extended to other domains involving Heterogeneous Information Networks (HINs), such as social networks and citation networks.
+  • We propose ProvSyn, a novel hybrid provenance graph synthesis framework that jointly models structural topology via graph generation models and textual semantics via fine-tuned LLMs, enabling the construction of high-fidelity provenance graphs.
+  • We introduce a novel five-dimensional fidelity evaluation framework specifically designed for provenance graph synthesis, encompassing structural, textual , temporal, embedding, and semantic metrics, addressing a critical gap in existing graph generation.
+  • Through systematic experiments on six benchmark datasets, we demonstrate that ProvSyn effectively mitigates data imbalance and enhances the generalizability of downstream apt detection models.
+  • We open source ProvSyn at https://anonymous.4open.science/r/OpenProvSyn-4D0D/ to facilitate future research in provenance graph synthesis and pids.
 2 Background and Motivation
-2.1 Class imbalance in Provenance Graphs
-We statistically analyzed the distributions of entity type and event type across four commonly-used public provenance datasets (in Appendix A Figure 3).
-Among the datasets, Cadets, Theia, and Trace are collected from DARPA Engagement 3, and Nodlink [37] is simulated provenance data collected on an Ubuntu system using Sysdig.
-The distributions exhibit a typical long-tail pattern and reveal significant data imbalance.
-In both entity type and event type, a single category accounts for over 50% of instances, exceeding 70% in some cases.
-In contrast, most other categories contribute less than 20%, with many below 5%.
-This imbalance further reflects the uneven distribution of system behaviors and attack patterns in most provenance datasets.
-The class imbalance problem poses a significant challenge for learning-based intrusion detection [21, 62, 74, 22, 39].
-Previous studies in the deep learning community have revealed that class imbalance has a detrimental effect on a model’s classification performance as well as its generalization abilities [8, 34].
-When deployed in open-world out-of-distribution (OOD) environments, detection models trained on imbalanced datasets may significantly exhibit performance degradation [67, 25].
-Hence, a more balanced dataset is vital for deep learning-based provenance-based IDS (PIDS) to achieve optimal results.
-However, this problem is largely ignored by the community, as there are very few studies considering this issue.
-2.2 Problems with Existing Methods
-To mitigate class imbalance, conventional techniques such as undersampling [44], oversampling [47], and the Synthetic Minority Over-Sampling Technique (SMOTE)  [9] have been proposed.
-However, these methods are not suitable for provenance graph data.
-Undersampling breaks the causality chain between system entities, while oversampling does not add variety to benign activities and can cause overfitting.
-SMOTE, which relies on linear interpolation in the feature space, cannot be applied to provenance graphs due to overlapping distributions of different classes in the feature space.
-Another approach for augmenting provenance datasets is to leverage graph generation models to synthesize provenance graphs.
-However, most existing models focus primarily on generating the structural connectivity among nodes and edges, without supporting additional node or edge attributes [12, 71].
-In contrast, provenance graphs contain rich textual attributes associated with nodes and edges, and the text attributes are crucial features for downstream intrusion detection systems [37, 53].
-Yet, to date, no graph generation model inherently supports the joint generation of both graph structure and textual attributes.
-With the recent advances in large language models, their capabilities in natural language understanding and generation have been widely demonstrated [4, 1].
-This motivates the exploration of the use of LLMs to synthesiz provenance data.
-Specifically, LLMs can be tasked with creating new provenance log entries.
-These generated logs contain rich texual and semantic information.
-However, when converting LLM-generated logs into graphs and evaluating them using graph-based metrics, their structural fidelity falls significantly short of graphs produced by heterogeneous graph generation models (see details in Section 5.2).
-This indicates that although current LLMs perform well in text synthesis, they lack a strong capability for synthesizing accurate graph topology and structure.
-To address these limitations, we propose ProvSyn, a framework that integrates the complementary strengths of graph generation models for structural synthesis and large language models for text attribute generation.
-3 Related Work
-Provenance-based Intrusion Detection System Provenance-based detection methods can be divided into three main categories [74]: rule-based, statistics-based, and learning-based techniques.
-Rule-based methods [24, 46, 28] rely on predefined heuristics derived from known attack patterns to identify malicious activities.
-Statistics-based [23, 61, 45] approaches analyze deviations in provenance graph elements by measuring their statistical anomalies.
-Learning-based techniques employ deep learning models to capture either normal system behavior [21, 62] or malicious patterns [41, 74, 39], framing APT detection as either a classification [39] or anomaly detection task [74, 22].
-These learning-based methods further branch into sequence-based approaches [41], which focus on execution workflows, and graph-based [21, 62, 74, 22, 39] methods, which use graph neural networks to model entity relationships and detect behavioral anomalies.
-Network Traffic Augmentation To address class imbalance in network traffic datasets, one line of work applies sampling techniques.
-Jiang et al.
-[32] adopt one-side selection to remove noisy majority samples and apply SMOTE [9] to augment minority instances.
-Liu et al.
-[42] propose the Difficult Set Sampling Technique to improve sample representativeness.
-Another line of research explores synthetic network traffic generation using deep learning.
-Generative models, especially GANs, have been employed for various scenarios, including flow-based traffic [55], manipulated traffic [11], and social media traffic [54], addressing challenges in realistic traffic synthesis [5].
-LLM-driven Synthetic Data Generation LLM-based synthetic data generation methods can be categorized into three main types: prompt engineering-based approaches [16, 20, 57, 43, 70, 38, 68], which utilize task descriptions, condition-value prompts, and in-context examples to steer generation; multi-step generation-based approaches [64, 15, 14, 63, 27], which break down complex tasks either at the sample or dataset level; and knowledge-enhanced approaches [50, 10, 30], which incorporate external knowledge such as knowledge graphs or web sources to improve the factual quality of outputs.
-The evaluation of synthetic data includes both direct metrics, which measure faithfulness [36] and diversity [73], and indirect metrics, which assess performance on downstream tasks through benchmark testing [58] and open-ended evaluation using human [26] or model-based [65] methods.
+2.1 Provenance Graph
+Provenance graphs are constructed from system audit events that record interactions among system entities, such as processes, files, and network connections.
+Formally, a provenance graph is a directed heterogeneous graph with textual attributes $G=(V,E)$, where each node $v_{i}\in V$ is associated with a type $t_{i}$ and a textual attribute: a name $n_{i}$, and each edge $e_{i}\in E$ is associated with a type $l_{i}$.
+Each system audit event is represented as a labeled 5-tuple: $e_{i}=(t_{i},n_{i},t_{j},n_{j},l_{i})$, where $t_{i}$ and $t_{j}$ denote the types of the source and target nodes, $n_{i}$ and $n_{j}$ denote their names, and $l_{i}$ denotes the edge type.
+Edges are directed from the source node to the target node, reflecting the causal relationships between entities.
+For node types, the most common node types include process, file, and network.
+For node names, process nodes are typically command lines, file nodes are their absolute file paths, and network nodes are their IP addresses.
+For edge types, common process–process interactions include CLONE, SENDMSG, and RECVMSG, which capture process creation and message-based communication.
+Common process–file interactions include READ, WRITE, and OPEN, which represent file access and modification behaviors.
+Common process–network interactions include CONNECT, SENDTO, and RECVFROM, which describe network connection establishment and data transmission.
+2.2 Class Imbalance
+Class imbalance poses a significant challenge in training effective machine learning models across diverse tasks, including financial fraud detection [dal2015calibrating, roy2024frauddiffuse], hate speech identification [davidson2017automated, caselli2021hatebert], and rare disease diagnosis [esteva2017dermatologist, tiu2022expert].
+Class imbalance refers to a dataset characteristic where the number of samples across classes is uneven, often with one or more classes dominating the dataset while the remaining classes account for only a small fraction [japkowicz2002class].
+Previous studies have revealed that class imbalance has a detrimental effect on a model’s classification performance as well as its generalization abilities in various machine learning tasks [buda2018systematic, johnson2019survey].
+When deployed in open-world out-of-distribution (OOD) environments, detection models trained on imbalanced datasets may significantly exhibit performance degradation [yang2024generalized, he2009learning].
+Hence, bridging the class imbalance gap in existing training datasets is vital for machine learning models to achieve robust results.
+Figure 1: Entity Type and Event Type Distribution in Provenance Dataset in OPTC-H501.
+To further illustrate the class imbalance issue in provenance datasets, we statistically analyze the distributions of entity types and event types across six commonly used public provenance datasets, including Cadets-E3, Theia-E3, Clearscope-E5, Theia-E5, OpTC-H201, and OpTC-H501.
+Figure 1 presents the distribution of OpTC-H501, while the distributions of the other datasets exhibit comparable distributional patterns and are shown in Appendix A (Figure 4).
+These distributions exhibit a typical long-tail pattern and reveal substantial data imbalance.
+For both entity types and event types, a single category accounts for more than 50% of the instances, exceeding 80% in some cases.
+In contrast, most other categories contribute less than 5%.
+This imbalance further reflects the uneven distribution of system behaviors in most provenance datasets.
+2.3 Limitations of Existing Methods
+To mitigate class imbalance, conventional techniques such as undersampling [liu2008exploratory], oversampling [mohammed2020machine], and the Synthetic Minority Over-Sampling Technique (SMOTE) [chawla2002smote] have been widely adopted.
+However, these methods are unsuitable for provenance graphs.
+Undersampling discards majority-class samples, disrupting causality chains between system entities.
+Oversampling replicates minority-class instances without introducing diversity, potentially causing overfitting.
+SMOTE generates synthetic samples through linear interpolation between neighbors in feature space; however, due to overlapping class distributions, it cannot be effectively applied to provenance graphs where structural validity must be preserved.
+An alternative approach leverages graph generation models to synthesize provenance graphs by learning underlying topology distributions.
+While frameworks such as MolGAN [de2018molgan], GraphRNN [you2018graph], and GraphGen [goyal2020graphgen] effectively capture structural connectivity, they primarily optimize adjacency patterns and categorical labels while neglecting textual attributes.
+In contrast, provenance graphs contain rich textual attributes associated with nodes—such as command lines, file paths, and IP addresses—that are crucial for downstream intrusion detection systems [li2023nodlink, rehman2024flash].
+To date, no graph generation model inherently supports joint generation of both graph structure and textual attributes.
+Recent advances in LLMs demonstrate strong capabilities in natural language understanding and generation [achiam2023gpt, claude], motivating exploration of LLM-based provenance data synthesis.
+Specifically, LLMs can generate provenance log entries containing rich textual and semantic information.
+However, when converting LLM-generated logs into graphs and evaluating them using graph-based metrics, their structural fidelity falls significantly short of graphs produced by heterogeneous graph generation models (see Section 5.2).
+This indicates that although current LLMs perform well in text synthesis, they lack strong capability for synthesizing accurate graph topology and structure.
+Figure 2: ProvSyn Architecture.
+First, a heterogeneous graph generation model constructs the initial structure, which is subsequently refined through rule-based topological constraints.
+Next, an LLM is fine-tuned on serialized provenance data to capture domain-specific semantics, enabling it to populate the generated nodes with synthesized textual attributes.
+Finally, the resulting graphs are validated via a fidelity evaluation framework across five distinct dimensions.
+3 Threat Model and Design Goal
+Threat Model.
+We follow the same threat model as past pids work [li2023nodlink, rehman2024flash, jia2024magic, jiang2025orthrus].
+We assume attackers originate externally and aim to compromise internal systems for valuable data or persistent control.
+Their actions, though potentially complex, are assumed to leave observable traces in system logs monitored by kernel-level tools.
+The system hardware, OS, and core auditing software form our trusted base.
+Poison attacks and evasion attacks are not considered in our threat model.
+Design Goal.
+ProvSyn is designed as a framework that integrates the complementary strengths of graph generation models for structural synthesis and large language models for text attribute generation.
+The objectives of ProvSyn are threefold: (1) to synthesize high-fidelity provenance graphs.
+High-fidelity provenance graphs require accurate modeling of both structural patterns and texual attributes; (2) to mitigate data imbalance in existing provenance graph datasets.
+ProvSyn is designed to learn the underlying distribution of real-world data and generate samples that represent underrepresented scenarios, thereby improving data diversity; and (3) to enhance the performance of downstream pids.
+By augmenting existing datasets with ProvSyn-generated graphs, the coverage of system behaviors can be expanded, leading to improved generalization and robustness of pids models.
 4 ProvSynDesign
-We implement the generation of provenance graphs in three stages.
-In the first stage, we train a heterogeneous graph generation model using real provenance graph dataset to produce the graph structure.
-After that, we apply rule-based post-processing to remove invalid edges and isolated nodes.
-In the third stage, we fine-tune a large language model to label the nodes with appropriate names.
-4.1 Problem Definition
-A system audit log dataset can be represented as a provenance graph, where nodes correspond to system entities and directed edges represent system events.
-Formally, a provenance graph is a directed heterogeneous graph $G=(V,E)$, where each node $v_{i}\in V$ is associated with a type $t_{i}=f(v_{i})$ and a name $n_{i}=h(v_{i})$, and each edge $e_{i}\in E$ is associated with a type $l_{i}=g(e_{i})$.
-Each system event is represented as a labeled 5-tuple:
-$e_{i}=(t_{i},n_{i},t_{j},n_{j},l_{i})$
-where $t_{i}$ and $t_{j}$ denote the types of the source and target nodes, $n_{i}$ and $n_{j}$ denote their names, and $l_{i}$ denotes the edge type.
-ProvSyn is to learn a distribution over such provenance graphs from real system audit data, enabling the generation of synthetic graphs that preserve both the structural and semantic properties of the original dataset.
-The generated graphs are also heterogeneous, with diverse node and edge types, and each node is assigned a detailed name.
-The objectives of ProvSyn are threefold: (1) to synthesize high-fidelity provenance graphs.
-This is our primary goal, as high-quality provenance data is essential for training reliable detection models; (2) to mitigate data imbalance in existing provenance graph datasets.
-ProvSyn is designed to learn the underlying distribution of real-world data and generate samples that represent underrepresented scenarios, thereby improving data diversity; and (3) to enhance the performance of downstream intrusion detection systems.
-By augmenting existing datasets with ProvSyn-generated graphs, the coverage of system behaviors can be expanded, leading to improved generalization and robustness of IDS models.
-Figure 1: ProvSyn Architecture.
-First, a heterogeneous graph generation model is employed to generate the initial structure of a provenance graph.
-Subsequently, the topology of the graph is refined according to predefined rules.
-Finally, a large language model is used to synthesize the textual attributes of the nodes within the provenance graph.
-4.2 Real-World Provenance Graph Construction
-We begin by parsing each log entry to extract five key elements: source node type, source node name, destination node type, destination node name, and edge type.
-These elements are used to construct a directed provenance graph, where each node is assigned two attributes (type and name), and each edge is labeled with a type attribute.
-To avoid duplication, we maintain a hash table indexed by the MD5 hash of node names.
-For log formats with simple key-value structures, we extract these elements directly using predefined keys.
-For formats with more complex attribute representations of entities and interactions, we apply regular expressions to extract the target elements.
-Extracting node names is particularly challenging; to improve accuracy, we define type-specific regular expressions tailored to each node type, ensuring that a valid name is extracted for every node.
-4.3 Graph Structure Generation
-Existing graph generation models primarily focus on structural connectivity, often overlooking the modeling of node and edge labels [12, 71].
-Many approaches also depend on domain-specific heuristics (e.g., those tailored to molecular graphs), limiting their generalizability across graph types [56].
-In contrast, provenance graphs represent a new class of heterogeneous graphs, where nodes correspond to system entities such as processes, files, and network components, and edges represent relations such as read, write, clone, and send.
+In this section, we introduce the design details of ProvSyn.
+Figure 2 shows the workflow of ProvSyn, which consists of three main components: graph structure synthesis, textual attribute synthesis, and fidelity evaluation framework.
+In the first stage, we train a heterogeneous graph generation model using real provenance graph dataset to generate the graph structure.
+In the second stage, we fine-tune a large language model to label the nodes with appropriate text attributes.
+Finally, we evaluate the fidelity of the synthesized provenance graphs from five perspectives.
+4.1 Graph Structure Synthesis
+In the first stage, we focus on generating the structure of provenance graphs.
+We first train a graph generation model on real provenance graph datasets, introduing the subgraph sampling strategy to address challenge of the scalability issue of large provenance graphs.
+Then for the synthesized graphs, we apply domain-specific refinement rules to ensure the validity of the generated graph structures.
+Structure Generation Model.
+Existing graph generation models primarily focus on structural connectivity, often overlooking the modeling of node and edge labels [de2018molgan, you2018graph].
+Many approaches also depend on domain-specific heuristics (e.g., those tailored to molecular graphs), limiting their generalizability across graph types [samanta2020nevae].
+In contrast, provenance graphs represent a new class of heterogeneous graphs, where nodes correspond to system entities and edges represent system call types.
 Generating such graphs requires a domain-agnostic model that captures both structural and label information.
-To this end, we adopt GraphGen [17], a domain-independent graph generation framework capable of jointly modeling node and edge labels along with structural connectivity.
-GraphGen converts graphs into sequences using minimum DFS codes, a canonical representation that preserves both topology and semantics.
-An Long Short-Term Memory (LSTM)-based model is trained to learn the joint distribution over these sequences, enabling the generation of labeled heterogeneous graphs in a domain-agnostic fashion.
-However, GraphGen is typically trained on small graphs with around 100 nodes, whereas provenance graphs often exceed 10,000 nodes, making it infeasible to feed the entire graph into the model.
-To make the training process scalable to large graphs, we adopt a random walk sampling strategy to extract subgraphs, and then the subgraphs can be used to construct a GraphGen-compatible training dataset.
-During sampling, we perform a restart-based random walk.
-At each step, there is a probability of returning to the starting node; otherwise, the next node is uniformly selected from the current node’s neighbors.
-Visited nodes and their connecting edges are dynamically added to the subgraph.
-The walk terminates when either a predefined number of iterations is reached or the subgraph exceeds the maximum allowed number of nodes or edges.
-The number of walks per starting node is determined by its degree, calculated as $\texttt{factor}\times\sqrt{\texttt{degree}}$ , ensuring that high-degree nodes generate more subgraphs.
-After sampling, each subgraph is normalized by remapping node IDs and removing self-loops.
-Only subgraphs that meet the node and edge limits and remain connected are retained.
-The reason why we adopt restart-based random walk for subgraph sampling is its ability to balance local exploration with broader graph coverage.
-Unlike simple random walk, which tends to get trapped in densely connected areas, and Metropolis-Hastings random walk, which often mixes slowly and introduces sampling bias in heterogeneous graphs, restart-based random walk periodically returns to the starting node.
-This mechanism helps it escape local regions and sample subgraphs that better reflect the overall graph structure.
-The complete procedure is presented in Algorithm 1.
-The training process of GraphGen involves converting graphs into minimum DFS codes, followed by learning the distribution of these sequences using a LSTM network.
-Given a graph dataset $\mathcal{G}=\{G_{1},G_{2},\dots,G_{m}\}$, each graph $G_{i}$ is first transformed into its minimum DFS code sequence $S_{i}=F(G_{i})=[s_{1},s_{2},\dots,s_{|E_{i}|}]$, where $|E_{i}|$ denotes the number of edges in graph $G_{i}$, and each $s_{t}$ is an edge tuple representing the $t$-th edge in the DFS traversal.
-Each edge tuple $s_{t}$ is defined as:
-$s_{t}=(t_{u},t_{v},L_{u},L_{e},L_{v})$
-where $t_{u}$ and $t_{v}$ are the timestamps (DFS indices) of the source and target nodes, respectively; $L_{u}$ and $L_{v}$ are the labels of the source and target nodes, respectively; and $L_{e}$ is the label of the edge between the two nodes.
-At each time step $t$, the LSTM updates its hidden state $h_{t}$ based on the embedding of the previous edge tuple $s_{t-1}$, and then predicts the distribution over the next edge tuple:
-$h_{t}=f_{\text{trans}}(h_{t-1},f_{\text{emb}}(s_{t-1}))$
-$s_{t}\sim p(s_{t}\mid h_{t})$
-Here, $f_{\text{emb}}$ is the embedding function that maps edge tuples to a vector space, and $f_{\text{trans}}$ is the transition function implemented by the LSTM.
-The training objective is to minimize the binary cross-entropy (BCE) loss between the predicted and ground-truth edge tuples across the entire dataset:
-$\mathcal{L}=\sum_{i=1}^{m}\sum_{t=1}^{|E_{i}|}\text{BCE}(p(s_{t}\mid s_{<t}),s_{t})$
-During inference, GraphGen generates new graphs using the trained LSTM model.
-Starting from an initial hidden state $h_{0}$ and a start-of-sequence symbol SOS, the model iteratively generates a minimum DFS code sequence $S=[s_{1},s_{2},\dots,s_{|E|}]$.
-The process of updating the hidden state and sampling the next edge tuple $s_{t}$ is the same as in the training phase.
-The generation continues until an end-of-sequence symbol EOS is produced.
-The resulting DFS code sequence $S$ is then converted back into a graph $G$ by reconstructing the edges based on the timestamps and labels in the sequence, yielding the generated graph structure and labels.
-This approach allows the model to generate graphs that reflect the structural and semantic patterns present in the training data.
-4.4 Rule-based Post-processing
-After the first stage, we obtain an undirected heterogeneous graph, where both nodes and edges have distinct types.
-However, this graph is not directly suitable as a provenance graph structure for two main reasons.
-First, provenance graphs are directed by nature, whereas the generated graph is undirected.
-Second, despite being trained on a large dataset, the heterogeneous graph generation model occasionally produces invalid connections—such as a process-read-process relation—which do not occur in real-world scenarios.
-To ensure the fidelity of generated provenance graphs, we implement a rule-based post-processing pipeline.
-First, we convert the undirected graph into a directed graph by replacing each undirected edge $(u,v)$ with two directed edges, $(u,v)$ and $(v,u)$.
-We then remove edges violating domain-specific constraints, such as connections between incompatible entity types or semantically invalid interactions.
-Finally, we remove all isolated nodes resulting from edge removal, as such nodes do not appear in real-world provenance graphs.
-The domain-specific constraints applied during post-processing vary by dataset.
-For example, for the Nodlink dataset, the constraints are as follows: (1) all operations must originate from process nodes; (2) file operations must terminate at file nodes; (3) process operations must terminate at process nodes; and (4) network operations must terminate at either network or file nodes.
-For the Cadets, Theia, and Trace datasets, the constraints are: (1) non-read operations must originate from process nodes; and (2) read operations (e.g., read, receive, load) must terminate at process nodes.
-[…]
+To this end, we adopt GraphGen [goyal2020graphgen], a domain-independent graph generation framework capable of jointly modeling node and edge labels along with structural connectivity.
+The GraphGen training process transforms each graph $G_{i}$ in a dataset $\mathcal{G}$ into its minimum DFS code sequence $S_{i}=[s_{1},s_{2},\dots,s_{|E_{i}|}]$, where each edge tuple $s_{t}=(t_{u},t_{v},L_{u},L_{e},L_{v})$ captures the DFS timestamps, node labels, and edge label.
+An LSTM network then learns the distribution of these sequences: at each step $t$, the hidden state $h_{t}$ is updated via the LSTM transition using the embedding of the previous edge tuple, and the model predicts the next tuple $s_{t}\sim p(s_{t}\mid h_{t})$.
+Training minimizes the binary cross-entropy loss $\mathcal{L}=\sum_{i}\sum_{t}\text{BCE}(p(s_{t}\mid s_{<t}),s_{t})$.
+During inference, the model starts from an initial state and a start symbol, then autoregressively samples a DFS code sequence until an end symbol appears.
+This sequence is finally converted back into a graph $G$ by reconstructing edges and labels.
+Input : Graph $G=(V,E)$, starting node $u\in V$, iterations $T$, restart probability $\alpha=0.15$, node constraints $\mathit{min\_nodes},\mathit{max\_nodes}$, edge constraints $\mathit{min\_edges},\mathit{max\_edges}$
 Output : Valid subgraph $G_{\text{sampled}}$ or $\emptyset$ if invalid
@@
 […]
-[…]
+3 $\mathit{current}\leftarrow u$
 4
@@
 7  if $r<\alpha$ then
-[…]
+     $\mathit{current}\leftarrow u$
      // Restart mechanism
@@
 9 else
-[…]
+10     $v\leftarrow$ random neighbor of $\mathit{current}$ in $G$
 11     if $v\notin V(G_{\text{sampled}})$ then
@@
 13
-[…]
-[…]
+14    Add edge $(\mathit{current},v)$ to $G_{\text{sampled}}$ (copy attributes from $G$)
+15     $\mathit{current}\leftarrow v$
 16
-[…]
+17 if $|V(G_{\text{sampled}})|\geq\mathit{max\_nodes}$ then
      break
@@
 18
-[…]
+19 if $|E(G_{\text{sampled}})|\geq\mathit{max\_edges}$ then
      break
@@
 20
-21Remove self-loops from $G_{\text{sampled}}$
+21
+22
+23Remove self-loops from $G_{\text{sampled}}$
 24 Relabel nodes to consecutive integers
@@
 26 $e\leftarrow|E(G_{\text{sampled}})|$
-25
-[…]
+27
+28if $\neg\text{is\_connected}(G_{\text{sampled}})$ or $n<\mathit{min\_nodes}$ or $n>\mathit{max\_nodes}$ or $e<\mathit{min\_edges}$ or $e>\mathit{max\_edges}$ then
   return $\emptyset$
   // Subgraph is invalid
-27
+29
 return $G_{\text{sampled}}$
 // Return valid subgraph
-28
+30
 Algorithm 1 Restart-Based Random Walk Subgraph Sampling
-4.5 Textual Attribute Generation
-After the first two stages, we obtain a directed heterogeneous graph in which nodes of different types are connected by edges of different types, and all connections conform to predefined rules.
-However, this graph is not yet sufficient to serve as a complete provenance graph.
-In real-world scenarios, nodes typically carry detailed names, such as command lines or file paths.
-These names are essential, as many provenance-based intrusion detection methods rely on features extracted from node names using natural language processing techniques [53, 37].
-Therefore, assigning meaningful node names is crucial for supporting accurate detection.
-In the third stage, we leverage the natural language generation capabilities of large language models to assign names to nodes in provenance graphs.
+Subgraph Sampling.
+However, GraphGen is typically trained on small graphs with around hunderds nodes, whereas provenance graphs contains millions of nodes, making it infeasible to feed the entire graph into the model.
+To make the training process scalable to large graphs, we adopt a random walk sampling strategy to extract subgraphs, and then the subgraphs can be used to construct a GraphGen-compatible training dataset.
+During sampling, we perform a rwr (rwr) subgraph sampling.
+RWR dynamically constructs subgraphs by performing random walks that probabilistically restart at the starting node, with the number of walks scaled by node degree to prioritize high-connectivity regions, and only retains connected, normalized subgraphs within predefined size constraints (Algorithm 1).
+Compared to other sampling methods [leskovec2006sampling, ahmed2014graph], RWR has three advantages: (1) Computational efficiency.
+RWR operates incrementally and does not require the full graph or global metrics, thus enabling efficient computation for provenance graphs with millions of nodes.
+(2) Enhanced local exploration.
+Compared to simple random walks, which may be confined to dense local regions, RWR incorporates restarts to return to the starting node, thereby enabling exploration across multiple neighborhoods.
+Unlike Metropolis–Hastings walks, which converge slowly in networks with skewed degree distributions, RWR adaptively explores local structure without requiring long mixing times [tong2006fast].
+(3) Generative modeling compatibility.
+By sampling connected subgraphs from a seed node through restarted walks, RWR avoids disconnected fragments produced by node- or edge-based methods [leskovec2006sampling, ahmed2014graph], resulting in more coherent training instances.
+Moreover, the termination conditions of RWR can be defined based on subgraph size, ensuring compatibility with the size constraints of downstream generative models.
+Rule-based Refinement.
+After structural generation, we obtain an undirected heterogeneous graph, where both nodes and edges have distinct types.
+However, the generated graph cannot be directly used as the structure of a provenance graph due to two main issues.
+First, provenance graphs are directed, whereas the generated graph is undirected.
+Second, despite being trained on large dataset, the heterogeneous graph generation model occasionally produces invalid connections (e.g., process-clone-network) relation, which do not occur in real-world scenarios.
+To ensure the fidelity of generated provenance graphs, we design a rule-based refinement process to convert the generated graph into a valid directed provenance graph.
+First, we convert the undirected graph into a directed graph by replacing each undirected edge $(u,v)$ with two directed edges, $(u,v)$ and $(v,u)$.
+We then remove edges violating domain-specific constraints, such as connections between incompatible entity types or semantically invalid interactions.
+Finally, we remove all isolated nodes resulting from edge removal.
+The rules are dataset-specific and are derived from the structural regularities observed in the corresponding real-world provenance graphs.
+Specifically, we define a set of permissible event types for each ordered pair of node types based on the ground-truth graph, and any edges in the generated graphs whose event types fall outside the predefined schema are filtered out during refinement.
+The complete and detailed rules for Darpa E3/E5 and OpTC are presented in Appendix B (Table 5).
+4.2 Textual Attribute Synthesis
+After the first stage, we obtain a directed graph in which nodes of different types are connected by edges of different types, and all connections conform to predefined rules.
+In this stage, we focus on assigning attributes to the nodes in the generated graph to synthesize complete provenance graphs.
+We leverage the natural language generation capabilities of large language models to assign names to nodes in provenance graphs.
 This process involves two main challenges.
-First, prior studies suggest that LLMs are not inherently effective at understanding graph structures  [33].
+First, prior studies suggest that LLMs are not inherently effective at understanding graph structures  [wang2023can, guo2023gpt4graph, jin2024large].
 Thus, a key issue is how to transform the structure of a provenance graph into a format that is more interpretable by LLMs.
 Second, LLMs lack prior knowledge of typical node names in provenance graphs, and prompting them directly often results in overly simplistic outputs with limited diversity.
-To address the first challenge, prior work has shown that LLMs are less effective at processing graph-structured data but perform well on sequence-based inputs [69].
+To address the first challenge, prior work has shown that LLMs are less effective at processing graph-structured data but perform well on sequence-based inputs [ye2023language, zhao2023graphtext, fatemi2023talk].
 Motivated by this, we propose a method that converts directed heterogeneous graphs into sequences of node and edge types using depth-first search.
@@
 Specifically, to extract complete and non-overlapping DFS paths, we use all nodes with zero in-degree as starting points for DFS traversal.
-However, due to the presence of self-loops in the provenance graph, some paths may not be captured using this strategy alone.
-To address this, we additionally include any node involved in a self-loop as a starting point.
 In each extracted sequence, a node is represented as a pair $(\tau_{i},n_{i})$, where $\tau_{i}\in\mathcal{T}$ denotes the node type and $n_{i}\in\mathcal{N}$ the node name.
@@
 To generate realistic Q&A pairs, we adopt a masking strategy that preserves structural context while requiring the model to infer node names.
-During the name generation process, two types of sequence are inputted to the LLM.
+During the inference process, two types of sequence are inputted to the LLM.
 The first type consists of sequences in which all nodes are unnamed, serving as a cold start and requiring the LLM to infer node names based solely on the order of node and edge types.
@@
 This design encourages the model to learn a mapping $f:(\tau_{i},\{\tau_{j},e_{l}\}_{j=1}^{k})\rightarrow n_{i}$, while preserving structural consistency through fixed sequences of node and edge types.
-[…]
+$Q_{\text{full}}$ $=[(\tau_{1},\text{[null]})\xrightarrow{e_{1}}(\tau_{2},\text{[null]})\xrightarrow{e_{2}}\cdots\xrightarrow{e_{k-1}}(\tau_{k},\text{[null]})],$
 $A_{\text{full}}$ $=[(\tau_{1},n_{1})\xrightarrow{e_{1}}(\tau_{2},n_{2})\xrightarrow{e_{2}}\cdots\xrightarrow{e_{k-1}}(\tau_{k},n_{k})].$
@@
 This approach leverages the strengths of LLMs in sequence-based tasks while addressing the challenges of graph structure understanding, leading to more effective node name generation in provenance graphs.
-4.6 Semantic Correctness Validation
-Evaluating semantic correctness is essential for assessing the quality of synthetic provenance graphs, as it indicates whether the generated graphs accurately reflect the underlying logic and intent of real-world system behaviors.
+4.3 Fidelity Evaluation Framework
+To comprehensively assess the fidelity of synthetic provenance graphs, we propose a framework that evaluates synthesized graphs from complementary perspectives, including structure, text attribute, temporal sequence, embedding, and semantics.
+Collectively, these evaluation protocols quantify the extent to which synthetic samples can serve as reliable surrogates for empirical system activities.
+Structural Evaluation.
+To quantitatively evaluate the structural fidelity of generated graphs, we employ Maximum Mean Discrepancy (MMD)-based metrics [gretton2006kernel, thompson2022evaluation] that measure distributional divergence between generated graphs and real graphs.
+These metrics include: (1) degree MMD for local connectivity patterns, (2) orbit MMD for higher-order structural motifs, (3) edge label MMD for edge label distributions, and (4) node label–degree MMD for label–topology interactions.
+To bridge the characterization of graph features with distributional divergence measurement, we leverage kernel methods to project these heterogeneous graphs into comparable representations.
+We construct kernel similarity matrices using two distinct Gaussian kernel formulations.
+The Earth Mover’s Distance-optimized Gaussian kernel is applied to categorical distribution metrics including degree MMD, edge label MMD, and node label-degree MMD, while the Euclidean-based Gaussian kernel handles continuous feature spaces in orbit MMD.
+The MMD² values are computed via the unbiased estimator, where $P$ and $Q$ denote reference and generated graph distributions, $x,x^{\prime}$ are samples from $P$, and $y,y^{\prime}$ from $Q$:
+$\text{MMD}^{2}=\$$\mathbb{E}_{x,x^{\prime}\sim P}[k(x,x^{\prime})]+\mathbb{E}_{y,y^{\prime}\sim Q}[k(y,y^{\prime})]$
+$-2\mathbb{E}_{x\sim P,y\sim Q}[k(x,y)]$
+Textual Evaluation.
+To evaluate the fidelity and diversity of generated text attributes, we adopt two standard NLP metrics: (1) GLEU [mutton2007gleu], which evaluates fidelity through n-gram overlap and is effective for short sequences; and (2) Distinct-1 [li2016diversity], which measures lexical diversity by computing the proportion of unique unigrams in the generated text.
+To adapt GLEU for graph textual attributes, we employ a reference-based greedy matching approach [rus2012optimal, kusner2015word, zhang2019bertscore] that focuses on the highest match to avoid dilution from irrelevant references.
+For each node type $t$, we construct a reference corpus $\mathcal{R}_{t}=\{r_{1},r_{2},\ldots,r_{n}\}$ by aggregating all valid node names from the real-world provenance graph.
+For a set of generated names $\mathcal{G}_{t}=\{g_{1},g_{2},\ldots,g_{|\mathcal{G}_{t}|}\}$, we compute similarities between each $g_{i}$ and all $r_{j}\in\mathcal{R}_{t}$, select the maximum similarity via greedy matching, and average these values:
+$S_{t}=\frac{1}{|\mathcal{G}_{t}|}\sum_{g_{i}\in\mathcal{G}_{t}}\max_{r_{j}\in\mathcal{R}_{t}}\text{GLEU}(g_{i},r_{j}).$
+We didn’t evaluate the text quality of network nodes because their names are purely numerical, making natural language metrics inappropriate.
+Additionally, our analysis shows that the network node names generated by both the baselines and ProvSyn conform to the IPv4 address format.
+Temporal Evaluation.
+A provenance graph is a directed graph in which directed sequences also represent temporal sequences of events occurring in chronological order.
+For synthetic provenance graphs, accurately preserving the temporal structure of event sequences is essential for ensuring fidelity.
+To evaluate the temporal consistency of synthetic provenance graphs, we employ two metrics: Longest Common Subsequence (LCS) [bergroth2000survey] and Dynamic Time Warping (DTW) [muller2007dynamic].
+LCS measures the longest event sequence appearing in both real and synthetic graphs in identical order.
+As a strict matching metric, LCS requires exact event alignment and temporal ordering without deviations.
+Conversely, DTW offers flexibility by allowing non-linear sequence alignments.
+Unlike strict methods, DTW enables sequences to be stretched or compressed along the time axis to accommodate local variations.
+This makes DTW suitable when synthetic graphs capture general patterns but exhibit slight temporal shifts.
+Embedding Evaluation.
+To evaluate the fidelity of synthetic graphs from embedding perspective, we employ graph representation learning techniques to project both synthetic and real graphs into a shared vector space, and measure their cosine similarity.
+Specifically, we adopt two representative embedding methods: NetLSD [tsitsulin2018netlsd] and Graph2Vec [narayanan2017graph2vec].
+NetLSD creates graph-level embeddings by summarizing a graph’s multi-scale structural properties using the heat kernel trace across different time scales.
+Its embedding vector is: $\mathbf{v}_{\text{netlsd}}=[\mathrm{tr}(e^{-t_{1}\mathbf{L}}),\mathrm{tr}(e^{-t_{2}\mathbf{L}}),\ldots,\mathrm{tr}(e^{-t_{K}\mathbf{L}})]$, where $\mathbf{L}$ is the graph’s Laplacian matrix, $t_{k}$ are diffusion time scales, and $\mathrm{tr}(\cdot)$ computes the matrix trace.
+Unlike NetLSD, Graph2Vec leverages node text attributes to model semantics and learns embeddings by modeling each graph as a corpus of WL-relabeled random walk sequences and applying document embedding techniques.
+Its embedding is defined as: $\mathbf{v}_{graph2vec}=\text{doc2vec}(\mathcal{W})$, where $\mathcal{W}=\{w_{1},w_{2},\ldots,w_{N}\}$ represents the set of random walk sequences generated from the WL-refined graph, and the doc2vec model produces the final graph embedding.
+To assess the similarity between a synthetic graph $G_{s}$ and a real graph $G_{r}$, we compute the cosine similarity between their respective embeddings:
+$\text{Similarity}=\frac{\mathbf{v}_{G_{s}}\cdot\mathbf{v}_{G_{r}}}{\|\mathbf{v}_{G_{s}}\|\|\mathbf{v}_{G_{r}}\|}$
+A higher similarity indicates that the synthetic graph exhibits greater fidelity to the real graph in the embedding space.
+Semantic Evaluation.
+Evaluating semantic correctness is essential for assessing the fidelity of synthetic provenance graphs, as it indicates whether the generated graphs accurately reflect the underlying logic and intent of real-world system behaviors.
 This evaluation examines whether the relationships among entities and the sequences of interactions conform to meaningful and valid operational patterns.
 Despite its importance, evaluating semantic correctness remains difficult due to the lack of standardized metrics and the inherent complexity in defining ground-truth semantics for diverse and dynamic system activities.
-Model-based Evaluation Mechanism.
 To evaluate semantic correctness, we formally define a semantic unit in the graph as a triplet: the starting node name $s$, edge type $e$, and ending node name $t$.
@@
 This sentence captures the semantics of the interaction between entities, expressing that the source node $s$ performs an action of type $e$ on the target node $t$.
-To assess the semantic correctness of $\text{Sentence}(s,e,t)$, we employ a BERT(Bidirectional Encoder Representations from Transformers) [13]-based encoder to obtain its embedding.
-This embedding captures the semantic representation of the interaction described by the triplet and serves as the basis for subsequent classification-based evaluation.
+To assess the semantic correctness of $\text{Sentence}(s,e,t)$, we employ a BERT(Bidirectional Encoder Representations from Transformers) [devlin2019bertpretrainingdeepbidirectional]-based encoder to obtain its embedding.
+This embedding captures the semantic representation of the interaction described by the triplet and serves as the basis for subsequent evaluation.
 Formally, let $\text{BERT}:\mathcal{S}\to\mathbb{R}^{d}$ be the encoding function that maps a sentence $\text{Sentence}(s,e,t)\in\mathcal{S}$ (the set of all possible semantic triplets in sentence form) to a dense vector $\mathbf{v}\in\mathbb{R}^{d}$.
-The embedding is computed as:
-$\mathbf{v}=\text{BERT}(\text{Sentence}(s,e,t))$
+The embedding is computed as $\mathbf{v}=\text{BERT}(\text{Sentence}(s,e,t))$.
 However, embedding of $\text{Sentence}(s,e,t)$ captures only pairwise semantic relationships, resulting in a localized evaluation.
-To achieve graph-level assessment, we employ a multi-layer Graph Attention Network (GAT) [60].
+To achieve graph-level assessment, we employ a multi-layer Graph Attention Network (GAT) [velickovic2017graph].
 This architecture aggregates neighborhood information through stacked self-attention layers, where each layer updates node representations by adaptively weighting features from adjacent nodes.
 We employ BERT to represent the text attributes of individual nodes as the initial node embeddings.
-Formally, for each node $v$ with text attribute $T_{v}$, its initial embedding $\mathbf{h}_{v}^{(0)}$ is derived via BERT:
-$\mathbf{h}_{v}^{(0)}=\text{BERT}(T_{v})$
+Formally, for each node $v$ with node name $T_{v}$, its initial embedding $\mathbf{h}_{v}^{(0)}$ is derived via BERT: $\mathbf{h}_{v}^{(0)}=\text{BERT}(T_{v})$.
 Let $\text{GAT}:\mathcal{V}\to\mathbb{R}^{d^{\prime}}$ denote the encoder that maps a node to its global representation in $\mathbb{R}^{d^{\prime}}$ using these initial embeddings.
@@
 $\mathbf{u}_{s}=\text{GAT}\left(s;\left\{\mathbf{h}_{v}^{(0)}\right\}_{v\in\mathcal{V}}\right),\quad\mathbf{u}_{t}=\text{GAT}\left(t;\left\{\mathbf{h}_{v}^{(0)}\right\}_{v\in\mathcal{V}}\right)$
-We then concatenate the global embeddings $\mathbf{u}_{s}$ and $\mathbf{u}_{t}$ with the local embedding $\mathbf{v}$ of the semantic unit to form a comprehensive representation:
-$\mathbf{z}=[\mathbf{v};\mathbf{u}_{s};\mathbf{u}_{t}]$
-This comprehensive representation $\mathbf{z}$ captures both the local and global context of the semantic unit.
-Finally, we feed $\mathbf{z}$ into a multilayer perceptron (MLP) discriminator to determine the correctness of the semantic unit.
-The MLP discriminator is defined as $\text{MLP}:\mathbf{R}^{d+2d^{\prime}}\to[0,1]$.
-The correctness score of the semantic unit is computed as:
-Correctness $\text{Correctness Score}=\text{MLP}(\mathbf{z})$
-The correctness score is a value between 0 and 1, where a higher score indicates higher semantic correctness.
-Contrastive Learning-based Training Framework.
-To train the GAT and MLP discriminator, we employ a contrastive learning approach.
-We use real provenance graphs as the training set, extracting all semantic units from them and labeling them as positive samples.
-Additionally, to provide negative samples for contrastive learning, we construct them from the positive samples using three distinct strategies to help the classifier model learn to distinguish between semantically correct and incorrect instances.
-The strategies are:
-Subject-Object Inversion (SOI): We swap the names of the start node and the end node.
-This trains the model to recognize cases where the subject and object are inverted in semantics.
-Formally, for a semantic unit $(s,e,t)$, we generate a negative sample by inverting the subject and object:
-$(s,e,t)\rightarrow(t,e,s)$
-Predicate Replacement (PR): We replace the edge type with another type of edge.
-This trains the model to identify incorrect predicates in semantics.
-However, in this method, replacing the edge type with another type may still result in semantically correct cases.
-To guarantee the constructed semantic units are semantically incorrect, we manually classify edge types into disjoint categories.
-Let $\mathcal{E}=\bigcup_{i=1}^{k}\mathcal{E}_{i}$ be the partition of edge types into $k$ categories.
-When replacing an edge type $e\in\mathcal{E}_{i}$, we ensure the new edge type $e^{\prime}$ is from a different category, i.e., $e^{\prime}\in\mathcal{E}_{j}$ where $i\neq j$.
-For a semantic unit $(s,e,t)$, we generate a negative sample by replacing the predicate:
-$(s,e,t)\rightarrow(s,e^{\prime},t)\quad\text{where }e^{\prime}\in\mathcal{E}_{j}\text{ and }j\neq i$
-Entity Substitution (ES): We replace the name of either the start node or the end node with the name of another node.
-This trains the model to recognize cases where irrelevant entities appear in semantics.
-It is important to note that in this construction method, replacing one element of a semantic triple with another element may still result in correct semantics.
-This is because the replaced element and the new element may be similar.
-Therefore, to construct negative samples, we first cluster the node names using K-means into disjoint clusters.
-Let $\mathcal{V}=\bigcup_{i=1}^{m}\mathcal{V}_{i}$ be the partition of node names into $m$ clusters.
-When replacing a node name $s\in\mathcal{V}_{i}$ or $t\in\mathcal{V}_{i}$, we select a node $s^{\prime}\in\mathcal{V}_{j}$ or $t^{\prime}\in\mathcal{V}_{j}$ where $i\neq j$.
-For a semantic unit $(s,e,t)$, we generate negative samples by substituting either the subject or the object:
-$(s,e,t)$ $\rightarrow(s^{\prime},e,t)\quad\text{where }s^{\prime}\in\mathcal{V}_{j}\text{ and }j\neq i$
-$\text{or}\quad(s,e,t)$ $\rightarrow(s,e,t^{\prime})\quad\text{where }t^{\prime}\in\mathcal{V}_{j}\text{ and }j\neq i$
-To maintain a balanced training set, we generate one negative sample for each positive sample in the real graph.
-Specifically, we randomly select one of the three strategies (SOI, PR, ES) to construct the negative sample.
-This ensures that positive and negative samples each account for 50% of the training data.
-For the MLP discriminator, which determines the semantic correctness of each triple, we adopt a binary cross-entropy loss.
-Given a batch of semantic units $\{(s_{i},e_{i},t_{i})\}_{i=1}^{N}$ with corresponding labels $y_{i}\in\{0,1\}$, the loss function is defined as:
-$\mathcal{L}_{\text{MLP}}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log f_{\text{MLP}}(\mathbf{z}_{i})+(1-y_{i})\log\left(1-f_{\text{MLP}}(\mathbf{z}_{i})\right)\right]$
+We then concatenate the global embeddings $\mathbf{u}_{s}$ and $\mathbf{u}_{t}$ with the local embedding $\mathbf{v}$ of the semantic unit to form a comprehensive representation.
+This comprehensive representation $\mathbf{z}=[\mathbf{v};\mathbf{u}_{s};\mathbf{u}_{t}]$ captures both the local and global context of the semantic unit.
+To quantify semantic correctness, we construct a reference semantic corpus from the real-world provenance graph by extracting all valid semantic units $(s,e,t)$ and encoding them using the same procedure described above, yielding a set of reference representations $\mathcal{Z}_{\text{ref}}=\{\mathbf{z}_{1},\mathbf{z}_{2},\ldots,\mathbf{z}_{m}\}$.
+Given a set of generated semantic units $\mathcal{Z}_{\text{gen}}$, we evaluate semantic correctness using the same greedy matching strategy [rus2012optimal, kusner2015word, zhang2019bertscore] as in the textual evaluation.
+Specifically, for each generated representation $\mathbf{z}\in\mathcal{Z}_{\text{gen}}$, we compute its maximum similarity with the reference corpus and then average over all generated units:
+$S_{\text{sem}}=\frac{1}{|\mathcal{Z}_{\text{gen}}|}\sum_{\mathbf{z}\in\mathcal{Z}_{\text{gen}}}\max_{\mathbf{z}_{j}\in\mathcal{Z}_{\text{ref}}}\text{sim}(\mathbf{z},\mathbf{z}_{j}).$
 5 Evaluation
 To thoroughly evaluate ProvSyn, we address the following research questions.
-Experiments on provenance graph structure generation were conducted on a machine equipped with Intel Xeon Silver CPU, 128 GB RAM, NVIDIA RTX 3090 GPUs, and Ubuntu 22.04.4 LTS.
-Experiments on text attribute synthesizing were performed on a machine with Intel Xeon Platinum CPU, 512 GB RAM, NVIDIA RTX A6000 GPUs, and Ubuntu 22.04.3 LTS.
+Experiments were performed on a machine with Intel Xeon Platinum CPU, 512 GB RAM, NVIDIA RTX A6000 GPUs, and Ubuntu 22.04.3 LTS.
 The research questions are:
-RQ1: What is the fidelity of synthetic provenance graphs in terms of structural, textual, temporal, and embedding-based characteristics?
-RQ2: What is the semantic correctness of synthetic provenance graphs?
-RQ3: Can synthetic provenance graphs mitigate class imbalance in the dataset?
-RQ4: When synthetic provenance graphs are used for training set augmentation in downstream intrusion detection tasks, what is the performance of the detection task?
-RQ5: What are the time and resource costs of generating synthetic provenance graphs?
-Further aspects of our study, including hyperparameter analysis, ablation study are discussed extensively in the Appendix C D.
+RQ1: What is the fidelity of synthetic provenance graphs in terms of structural, textual, temporal, embedding, and semantic characteristics?
+RQ2: Can synthetic provenance graphs mitigate class imbalance in the dataset?
+RQ3: When synthetic provenance graphs are used for training set augmentation in downstream intrusion detection tasks, what is the performance of the detection task?
+RQ4: What are the time and resource costs of generating synthetic provenance graphs?
+Further aspects of our study, including hyperparameter analysis, ablation study are discussed in the Appendix D and  E.
 5.1 Experimental Setup
 Datasets.
-We evaluate the effectiveness of ProvSyn on two public datasets: the Nodlink Ubuntu dataset [37] and the DARPA Engagement 3 datasets [2].
-The Nodlink dataset [37] we use is a simulated provenance dataset collected on Ubuntu 20.04 using Sysdig.
-It provides both benign logs and attack logs containing a complete multi-stage attack chain.
-The detailed statistics of the Nodlink provenance graphs used for training and evaluation are shown in Table I.
-The DARPA E3 datasets [2] are collected from an enterprise network during adversarial engagements as part of the DARPA Transparent Computing program.
-APT attacks are launched by a red team, while blue teams attempt to detect them through host auditing and causality analysis.
-TRACE, THEIA and CADETS sub-datasets are included in our evaluation.
-For E3 dataset, we utilize labels from ThreaTrace.
-As shown in Table I, the E3 dataset, compared to Nodlink, has a larger number of nodes and edges, and also exhibits greater diversity in node and edge types.
-TABLE I: Comparison of dataset statistics.
-Dataset Nodes Edges Node Types Edge Types
-Nodlink 16,818 30,531 3 13
-Cadets 89,207 276,883 6 20
-Theia 77,071 202,517 3 14
-Trace 74,241 113,808 7 20
+We evaluate the effectiveness of ProvSyn on three public datasets: the DARPA Engagement 3 dataset [darpa-E3], the DARPA Engagement 5 dataset [darpa-E5], and the OpTC dataset [optc-data], utilizing the labels from Orthrus[jiang2025orthrus].
+From each of these datasets, we select two subsets, resulting in a total of six datasets: Cadets-E3, Theia-E3, Clearscope-E5, Theia-E5, OpTC-H501, and OpTC-H201.
+The provenance graphs in these datasets contain three types of nodes and ten types of events, with the number of nodes ranging from 50k to 2,000k and the number of edges ranging from 100k to 8,000k.
+For each dataset, we use provenance data from the first day for training and fidelity evaluation, with the two sets being derived from distinct time windows and strictly separated.
 Baselines.
@@
 This method does not require an additional heterogeneous graph generation model.
-We systematically evaluate and compare the provenance data synthesized by these strong LLMs with the data synthesized by ProvSyn.
-To ensure a comprehensive comparison, we selected baseline models encompassing both large-parameter closed-source models and smaller-parameter open-source models.
-For large-parameter LLMs, we chose GPT-4o [4], Claude-3.7-Sonnet [1], and Doubao-pro-32k [3].
-For smaller-parameter LLMs, we chose Qwen2.5-7B-Instruct [66] and Llama3.1-8B-Instruct [18].
-To generate provenance logs using LLMs listed above, we designed structured prompts that included task descriptions, illustrative examples, and clearly defined output formats.
+We systematically evaluate and compare the provenance data synthesized by LLMs with the data synthesized by ProvSyn.
+To ensure a competitive comparison, we selected a set of strong baseline models, including GPT-5.1 [openai-gpt51], Claude-Sonnet-4.5 [claude], DeepSeek-v3.2 [liu2025deepseek], and Qwen3-Max [yang2025qwen3].
+These models are widely regarded as state-of-the-art LLMs at the time of evaluation.
+To generate provenance logs using LLMs listed above, we designed structured prompts that included task descriptions and illustrative examples.
 Each dataset was paired with a customized prompt, primarily differing in the examples provided.
-These examples, formatted as JSON or TSV entries, contained five elements: source node type, source node name, destination node type, destination node name, and edge type.
+These examples contained five elements: source node type, source node name, destination node type, destination node name, and edge type.
 They were selected from real audit logs to ensure both representativeness and coverage of various node and edge types.
 For each dataset, we collected 100 responses from each LLM.
-After obtaining the responses, we parsed them line by line and filtered out entries that did not conform to the expected JSON or TSV format.
-We also discarded entries containing node or edge types outside the predefined type sets.
-The remaining entries were then used to construct provenance graphs for comparison with graphs generated by ProvSyn.
+After obtaining the responses, we filter out entries that contain node or edge types outside of the predefined type sets.
+The remaining entries were then used to construct graphs for comparison with graphs generated by ProvSyn.
+The prompt used in our experiments is provided in Appendix C (Figure 7).
+Table 1: Comprehensive Comparison of MMD, Temporal, and Embedding Metrics.
+For MMD and DTW, lower is better.
+For LCS, NetLSD, and Graph2Vec, higher is better.
+Bold values indicate the best performance in each category.
+Dataset Model MMD Metrics ($\downarrow$) Temporal Metrics Embedding Metrics ($\uparrow$)
+Degree Orbit Edge L. Node L.+Deg LCS ($\uparrow$) DTW ($\downarrow$) NetLSD Graph2Vec
+Cadets_E3 GPT-5.1 0.59 0.87 0.41 1.16 3.02 1.58 0.48 0.29
+Claude-sonnet-4.5 1.01 0.96 0.35 1.28 4.07 1.62 0.16 0.33
+DeepSeek-v3.2 1.47 0.88 0.72 1.45 3.03 1.69 0.03 0.34
+Qwen3-max 1.36 0.94 0.10 1.28 3.00 0.00 0.03 0.29
+ProvSyn 0.37 0.79 0.01 0.61 4.71 0.74 0.49 0.35
+Theia_E3 GPT-5.1 0.28 0.37 0.17 0.60 3.14 0.53 0.34 0.36
+Claude-sonnet-4.5 0.42 0.33 0.19 0.73 3.00 0.00 0.83 0.34
+DeepSeek-v3.2 0.57 0.38 0.29 0.65 2.93 1.92 0.42 0.38
+Qwen3-max 0.81 0.90 0.22 1.07 3.00 0.00 0.85 0.31
+ProvSyn 0.07 0.19 0.07 0.29 3.19 0.94 0.53 0.47
+ClearScope_E5 GPT-5.1 0.73 0.47 0.01 1.16 3.00 1.64 0.54 0.18
+Claude-sonnet-4.5 0.61 1.01 0.02 1.11 3.00 1.71 0.60 0.19
+DeepSeek-v3.2 0.56 0.49 0.01 0.93 3.00 1.64 0.41 0.19
+Qwen3-max 0.57 0.59 0.00 0.64 3.00 1.66 0.62 0.19
+ProvSyn 0.22 0.38 0.00 0.23 3.23 1.48 0.73 0.34
+Theia_E5 GPT-5.1 0.66 0.60 0.12 0.67 3.03 0.45 0.42 0.37
+Claude-sonnet-4.5 0.60 0.72 0.08 0.81 3.21 0.46 0.36 0.37
+DeepSeek-v3.2 0.70 1.10 0.43 0.54 3.02 0.53 0.10 0.40
+Qwen3-max 0.66 1.06 0.18 0.69 3.03 0.42 0.23 0.37
+ProvSyn 0.13 0.29 0.03 0.20 3.02 0.35 0.53 0.35
+OpTC_H201 GPT-5.1 0.65 0.75 0.51 0.62 4.09 2.98 0.28 0.23
+Claude-sonnet-4.5 0.98 1.22 1.93 1.77 3.50 3.04 0.04 0.21
+DeepSeek-v3.2 0.59 0.96 1.17 1.15 2.86 3.32 0.01 0.28
+Qwen3-max 0.55 0.79 1.00 1.05 2.76 3.36 0.01 0.29
+ProvSyn 0.10 0.51 0.11 0.17 4.74 1.73 0.64 0.33
+OpTC_H501 GPT-5.1 0.85 1.08 1.13 1.04 3.86 2.56 0.14 0.31
+Claude-sonnet-4.5 0.45 1.09 0.92 1.17 3.90 2.74 0.03 0.26
+DeepSeek-v3.2 0.77 1.19 0.93 1.52 2.45 3.71 0.01 0.37
+Qwen3-max 0.67 1.09 1.54 1.38 2.64 3.30 0.01 0.37
+ProvSyn 0.08 0.52 0.02 0.11 4.65 1.78 0.61 0.40
+Metrics.
+For fidelity evaluation, we adopt the five-dimensional metrics introduced in Section 4.3.
+For class imbalance mitigation, we use three metrics: novelty, uniqueness, and normalized entropy.
+Novelty verifies that each generated graph is neither a subgraph of any training reference graph nor contains any reference graph as a subgraph, indicating the presence of novel structures.
+Uniqueness measures the proportion of generated graphs that are structurally distinct from one another, reflecting diversity.
+Normalized entropy measures label uncertainty and rescales it by the maximum entropy under the same number of classes, yielding values in the range $[0,1]$.
+Higher values indicate a more balanced label distribution.
+For IDS application, beyond the standard F-score, we report three robust metrics for comprehensive evaluation: Balanced Accuracy (B.Acc), which reduces the impact of imbalanced test data; the Matthews Correlation Coefficient (MCC) [jiang2025orthrus], which measures the correlation between predicted and true labels; and the Diagnostic Odds Ratio (DOR), which reflects diagnostic performance as the odds ratio of positive results for attacks versus normal behaviors.
 Configurations.
-For training the heterogeneous graph generation model, we sampled 20,000 subgraphs from the real provenance graphs using a random walk strategy.
+For training the graph generation model, we sampled 20,000 subgraphs from the real provenance graphs using RWR subgraph sampling strategy.
 Among them, 16,000 subgraphs were used for training, 2,000 for validation, and 2000 for test.
 The model was trained for 3,000 epochs with a batch size of 32 and a learning rate of 0.003.
-After training, we selected the model that achieved the best performance on the validation set for use in ProvSyn.
-For example, on the Theia dataset, the best-performing model was obtained after 2,380 epochs and was used in the final ProvSyn pipeline.
-For the text attribute synthesizing task, we used Llama3.2-3B-Instruct as the base model.
+For the text attribute synthesizing task, we used Llama3.2-3B-Instruct [grattafiori2024llama] as the base model.
 We constructed a training set of 20,000 samples, consisting of 10,000 fully masked QA pairs and 10,000 partially masked QA pairs.
@@
 5.2 Fidelity Evaluation
-TABLE II: Comparison of MMD metrics across different models and datasets. Lower values indicate better performance.
-Dataset Model Degree Orbit Node Label Edge Label Node Label & Degree
-Nodlink GPT-4o 0.75 0.94 0.11 0.75 0.89
-Claude-3.7 0.56 0.85 0.07 0.92 0.77
-Doubao-pro 0.58 0.82 0.18 1.50 0.59
-Qwen2.5-7B 0.43 0.59 0.05 0.94 0.40
-Llama3.1-8B 0.62 - 0.10 0.91 0.81
-ProvSyn 0.07 0.07 0.002 0.01 0.07
-Cadets GPT-4o 0.82 0.28 0.37 0.90 0.48
-Claude-3.7 0.76 0.36 0.11 1.33 0.83
-Doubao-pro 0.73 0.25 0.27 1.31 0.66
-Qwen2.5-7B 0.76 0.39 0.67 0.77 0.33
-Llama3.1-8B 0.88 0.40 0.72 0.85 0.81
-ProvSyn 0.08 0.32 0.003 0.01 0.12
-Theia GPT-4o 0.94 0.35 0.46 1.34 0.53
-Claude-3.7 0.92 0.85 0.18 1.44 0.81
-Doubao-pro 0.82 0.67 0.46 1.86 0.89
-Qwen2.5-7B 0.91 0.46 0.48 1.00 0.61
-Llama3.1-8B 0.68 0.87 0.53 1.01 0.66
-ProvSyn 0.16 0.54 0.005 0.01 0.29
-Trace GPT-4o 0.66 1.04 0.80 1.07 0.54
-Claude-3.7 0.77 1.10 0.68 1.74 0.80
-Doubao-pro 0.75 1.06 0.57 1.12 0.51
-Qwen2.5-7B 0.24 0.23 0.44 1.06 0.47
-Llama3.1-8B 0.25 0.87 1.04 1.16 0.63
-ProvSyn 0.17 0.23 0.02 0.04 0.33
+Table 2: Comparison of textual quality across different models, node types, and datasets. Left: GLEU. Right: Distinct-1. The maximum value in each group is bolded.
+Dataset Type GLEU Distinct-1
+GPT Claude DeepSeek Qwen3 ProvSyn GPT Claude DeepSeek Qwen3 ProvSyn
+Cadets-E3 Process 0.36 0.16 0.04 0.07 0.56 1.00 1.00 1.00 1.00 1.00
+File 0.10 0.07 0.02 0.03 0.22 1.00 1.00 1.00 1.00 1.00
+Theia-E3 Process 0.03 0.14 0.00 0.50 0.69 0.66 1.00 0.51 1.00 0.70
+File 0.01 0.00 0.00 0.00 0.06 1.00 1.00 1.00 1.00 1.00
+ClearScope-E5 Process 0.47 0.78 0.37 0.33 0.52 0.55 0.67 0.51 0.57 0.68
+File 0.08 0.06 0.06 0.08 0.21 0.75 0.81 0.73 0.73 0.84
+Theia-E5 Process 0.04 0.05 0.03 0.05 0.18 0.54 0.60 0.38 0.57 0.83
+File 0.03 0.03 0.03 0.03 0.03 1.00 1.00 1.00 1.00 1.00
+OpTC-H201 Process 0.21 0.33 0.19 0.19 0.55 0.31 0.38 0.22 0.40 0.46
+File 0.17 0.15 0.15 0.19 0.30 0.70 0.94 0.95 0.99 0.71
+OpTC-H501 Process 0.33 0.36 0.23 0.31 0.38 0.20 0.25 0.34 0.31 0.39
+File 0.00 0.00 0.00 0.00 0.00 0.69 0.76 0.80 0.79 0.84
 Structural Evaluation.
-To quantitatively evaluate the structural fidelity of generated graphs, we employ five Maximum Mean Discrepancy (MMD)-based metrics [19, 59] that measure distributional divergence between generated graphs and real graphs.
-These metrics include: (1)degree MMD for local connectivity patterns, (2)orbit MMD for higher-order structural motifs, (3)node label MMD for vertex attribute distributions, (4)edge label MMD for edge label distributions, and (5)joint node label-degree MMD for attribute-topology interactions.
-To bridge the characterization of graph features with distributional divergence measurement, we leverage kernel methods to project these heterogeneous graphs into comparable representations.
-We construct kernel similarity matrices using two distinct Gaussian kernel formulations.
-The Earth Mover’s Distance-optimized Gaussian kernel is applied to categorical distribution metrics including degree MMD, node/edge label MMD, and joint node label-degree MMD, while the Euclidean-based Gaussian kernel handles continuous feature spaces in orbit MMD.
-The MMD² values are computed via the unbiased estimator:
-$\text{MMD}^{2}=$ $\mathbb{E}_{x,x^{\prime}\sim P}[k(x,x^{\prime})]+\mathbb{E}_{y,y^{\prime}\sim Q}[k(y,y^{\prime})]$
-$-2\mathbb{E}_{x\sim P,y\sim Q}[k(x,y)]$
-where $P$ and $Q$ denote reference and generated graph distributions, $x,x^{\prime}$ are samples from $P$, and $y,y^{\prime}$ from $Q$.
-In the experimental results shown in Table II, ProvSyn consistently outperformed the baselines across all datasets.
-In Nodlink dataset, ProvSyn achieved values an order of magnitude lower than the baselines on all metrics, indicating superior structural fidelity.
-In Cadets, Theia, and Trace datasets, ProvSyn also maintained a clear lead, particularly on label-related metrics.
-In contrast, baseline methods exhibit inconsistent performance patterns across the four datasets, suggesting fundamental limitations in structural generation capabilities regardless of the scale of model parameters.
-This finding underscores the necessity of an additional heterogeneous graph generation network, which produces structurally superior graphs compared to those generated solely by an LLM.
-Texutual Evaluation.
-To evaluate the accuracy of generated text attributes, we use standard NLP metrics that capture different types of similarity.
-These include: (1) BLEU [51], which measures n-gram precision; (2) GLEU [49], which balances precision and recall; and (3) ROUGE [40], which focuses on recall by assessing how much of the reference text is covered.
-For each node type, we construct a reference corpus $\mathcal{R}_{t}=\{r_{1},r_{2},...,r_{n}\}$ by aggregating all valid node names of type $t$ from the real-word provenance graph.
-When evaluating a generated name $g_{i}$, we compute its similarity with every reference $r_{j}\in\mathcal{R}_{t}$ through parallel metric calculations, then select the maximum similarity score as the final assessment:
-$S(g_{i})=\max_{r_{j}\in\mathcal{R}_{t}}\text{sim}(g_{i},r_{j})$
-The choice of maximum similarity over average similarity is driven by the following considerations.
-Focusing on the highest match ensures that the generated text is evaluated based on its optimal alignment with the reference corpus, thereby highlighting its capacity to produce accurate and relevant attributes.
-The complete procedure is presented in Algorithm 2 in Appendix B.
-In our experiments, we used the Python NLTK library to compute the BLEU, GLEU, and ROUGE scores for the process and file nodes.
-We didn’t evaluate the text quality of network nodes because their names are purely numerical, making natural language metrics inappropriate.
-Additionally, our analysis shows that the network node names generated by both the baselines and ProvSyn conform to the IPv4 address format.
-The results in Table III show that the node names generated by the baselines generally achieve scores below 0.1, indicating that large language models lack prior knowledge of the names of operating system entities.
-In contrast, ProvSyn was exposed to real entity names from audit logs during the finetuning stage, enabling a smaller 3B model to generate more accurate and realistic names.
-Interestingly, we observed that Llama-8B performed relatively well in the Cadets datasets.
-Further analysis revealed that many of the node names generated by Llama were directly copied from the examples in the prompt, even though the corresponding log entries did not make sense.
+Table 1 reports the structural fidelity between generated graphs and real data as measured by MMD metrics.
+Across all datasets, ProvSyn consistently achieves the lowest scores on degree, orbit, edge label, and node label &degree distributions, indicating closer alignment with the underlying graph structures.
+Notably, on the OpTC-H201 and OpTC-H501 datasets, ProvSyn reduces MMD scores by more than one order of magnitude compared to all baseline methods.
+In contrast, baseline models exhibit higher and less stable scores across datasets, suggesting limited ability to preserve structural fidelity.
+These results highlight the importance of incorporating a dedicated heterogeneous graph generation network to achieve faithful structural synthesis beyond standalone LLM capabilities.
+Textual Evaluation.
+In our experiments, we used the Python NLTK library to compute the GLEU, and Distinct-1 scores for the process and file nodes.
+The results in Table 2 show that the GLEU score of node names generated by the baselines generally achieve scores below 0.20, indicating that large language models lack prior knowledge of the names of operating system entities.
+In contrast, ProvSyn was exposed to real entity names during the finetuning stage, enabling a smaller model to generate more accurate and realistic names.
+Regarding the Distinct-1 diversity metric, ProvSyn outperforms baseline models across most datasets and node types.
+In particular, for process nodes in ClearScope-E5 and Theia-E5, baseline models exhibit lower Distinct-1 scores, reflecting repeated or monotonous name generation.
+By contrast, ProvSyn produces more distinctive and diverse entity names, better reflecting the complex naming distributions of operating systems.
 Temporal Evaluation.
-A provenance graph is a directed graph in which directed sequences also represent temporal sequences of events occurring in chronological order.
-For synthetic provenance graphs, accurately preserving the temporal structure of event sequences is essential for ensuring fidelity.
-To evaluate the temporal consistency of synthetic provenance graphs, we employ two metrics: Longest Common Subsequence (LCS) [7] and Dynamic Time Warping (DTW) [48].
-LCS measures the length of the longest sequence of events that appears in both the real and synthetic graphs in the same order.
-LCS is a strict matching metric—it requires exact alignment of events and their temporal order without allowing for any deviations or shifts.
-DTW, on the other hand, is a more flexible metric that allows for non-linear alignments between sequences.
-Unlike strict matching methods, DTW allows sequences to be stretched or compressed along the time axis, thereby accommodating local variations.
-This flexibility makes DTW particularly suitable for scenarios where the synthetic graph captures the general pattern of the real sequence but exhibits slight shifts or inconsistencies.
 In our experiments, we used depth-first search to extract sequences from the graphs.
@@
 For LCS computation, we used a dynamic programming approach, and for DTW, we used the dtw library in Python.
-To compare synthetic graphs with the real graphs in the test set, we adopted a strategy similar to Algorithm 2 in Appendix B: each sequence from the synthetic graph was compared with all sequences from the real graph, and the best match was recorded as its score.
-The overall graph score was then computed by aggregating the scores of all sequences in the synthetic graph.
-The results in Table IV show that ProvSyn achieves higher LCS scores and lower DTW values than the baselines across all four datasets, indicating its advantage in preserving temporal consistency.
-TABLE III: Comparison of textual quality across different models, types and datasets. Lower values indicate better performance.
-Dataset Type Model BLEU GLEU ROUGE-L-F
-Nodlink Process GPT-4o 0.05 0.11 0.23
-Claude-3.7 0.08 0.15 0.27
-Doubao-pro 0.02 0.03 0.09
-Qwen2.5-7B 0.07 0.15 0.29
-Llama3.1-8B 0.12 0.17 0.33
-ProvSyn 0.89 1.00 0.99
-File GPT-4o 0.09 0.11 0.29
-Claude-3.7 0.10 0.11 0.29
-Doubao-pro 0.06 0.11 0.25
-Qwen2.5-7B 0.11 0.17 0.41
-Llama3.1-8B 0.26 0.26 0.37
-ProvSyn 0.31 0.32 0.44
-Cadets Process GPT-4o 0.24 0.24 0.24
-Claude-3.7 0.17 0.17 0.17
-Doubao-pro 0.03 0.03 0.03
-Qwen2.5-7B 0.03 0.03 0.04
-Llama3.1-8B 1.00 1.00 1.00
-ProvSyn 1.00 1.00 1.00
-File GPT-4o 0.01 0.01 0.20
-Claude-3.7 0.03 0.03 0.14
-Doubao-pro 0.01 0.01 0.14
-Qwen2.5-7B 0.02 0.02 0.16
-Llama3.1-8B 0.00 0.00 0.03
-ProvSyn 0.31 0.32 0.60
-Theia Process GPT-4o 0.08 0.08 0.09
-Claude-3.7 0.04 0.04 0.05
-Doubao-pro 0.01 0.01 0.03
-Qwen2.5-7B 0.19 0.19 0.20
-Llama3.1-8B 0.00 0.00 0.00
-ProvSyn 0.76 0.89 0.47
-File GPT-4o 0.03 0.03 0.21
-Claude-3.7 0.05 0.05 0.16
-Doubao-pro 0.01 0.01 0.11
-Qwen2.5-7B 0.00 0.00 0.29
-Llama3.1-8B 0.00 0.00 0.35
-ProvSyn 0.37 0.44 0.48
-Trace Process GPT-4o 0.02 0.02 0.02
-Claude-3.7 0.02 0.02 0.02
-Doubao-pro 0.00 0.00 0.00
-Qwen2.5-7B 0.02 0.02 0.02
-Llama3.1-8B 0.00 0.00 0.00
-ProvSyn 0.50 0.50 0.50
-File GPT-4o 0.00 0.00 0.03
-Claude-3.7 0.00 0.00 0.03
-Doubao-pro 0.00 0.00 0.01
-Qwen2.5-7B 0.00 0.00 0.02
-Llama3.1-8B 0.00 0.00 0.00
-ProvSyn 0.00 0.00 0.67
-Graph Embedding Evaluation.
-To evaluate the fidelity of synthetic graphs from an embedding perspective, we employ graph representation learning techniques to project both synthetic and real graphs into a shared vector space, and measure their similarity using cosine similarity.
-Specifically, we adopt two representative embedding methods: DeepWalk [52] and Doc2Vec [35].
-DeepWalk is a foundational algorithm in graph representation learning.
-It learns node embeddings by simulating truncated random walks on the graph and treating the resulting node sequences as sentences for a language modeling task.
-Given a graph $G$, we perform $N$ random walks of length $L$ starting from each node, generating a corpus $\mathcal{W}=\{w_{1},w_{2},\ldots,w_{N}\}$, where each $w_{i}$ is a walk sequence of node identifiers.
-These sequences are then used to train a Word2Vec model, producing node-level embeddings.
-The graph-level embedding $\mathbf{v}_{\text{deepwalk}}$ is computed as the mean of all node embeddings:
-$\mathbf{v}_{\text{deepwalk}}=\frac{1}{|\mathcal{V}|}\sum_{v\in\mathcal{V}}\text{deepwalk}(v)$
-Notably, DeepWalk does not consider textual attributes of nodes.
-Unlike DeepWalk, Doc2Vec leverages node name information to capture the textual semantics of the graph.
-It treats each graph as a document, which is constructed from a collection of random walks.
-For a given graph $G$, we perform $N$ random walks of length $L$, resulting in a sequence set $\mathcal{W}=\{w_{1},w_{2},\ldots,w_{N}\}$, where each $w_{i}$ is a sequence of node names.
-These sequences are used to train a Doc2Vec model, yielding a graph-level embedding $\mathbf{v}_{\text{doc2vec}}$:
-$\mathbf{v}_{\text{doc2vec}}=\text{doc2vec}(\mathcal{W})$
-To assess the similarity between a synthetic graph $G_{s}$ and a real graph $G_{r}$, we compute the cosine similarity between their respective embeddings:
-$\text{Similarity}=\frac{\mathbf{v}_{G_{s}}\cdot\mathbf{v}_{G_{r}}}{\|\mathbf{v}_{G_{s}}\|\|\mathbf{v}_{G_{r}}\|}$
-A higher similarity score indicates that the synthetic graph exhibits greater fidelity to the real graph in the embedding space.
-For DeepWalk-based similarity, we use the Word2Vec model from the Gensim library, and for Doc2Vec-based similarity, we use Gensim’s Doc2Vec model.
-In our experiments results shown in Table IV, ProvSyn outperforms the baselines on the Nodlink, Cadets, and Trace datasets, while performing slightly worse than GPT and Claude on Theia.
-The results show that ProvSyn maintains stable embedding similarity, with DeepWalk scores above 0.80 and Doc2Vec scores above 0.65 across four datasets.
-In comparison, the baselines exhibit greater variation across different datasets.
-TABLE IV: Comparison of Temporal and Embedding metrics across different datasets and models. For LCS, higher is better; for DTW, lower is better; for DeepWalk and Doc2Vec, higher is better.
-Dataset Model Temporal Metrics Embedding Metrics
-LCS DTW DeepWalk Doc2Vec
-Nodlink GPT-4o 2.44 1.54 0.66 0.69
-Claude-3.7 3.03 1.33 0.83 0.40
-Doubao-pro 2.64 0.89 0.60 0.59
-Qwen2.5-7B 2.84 0.56 0.58 0.32
-Llama3.1-8B 2.88 0.83 0.50 0.16
-ProvSyn 3.30 0.41 0.89 0.74
-Cadets GPT-4o 2.99 0.48 0.70 0.40
-Claude-3.7 2.87 1.20 0.73 0.39
-Doubao-pro 2.68 3.62 0.64 0.71
-Qwen2.5-7B 2.13 24.30 0.79 0.25
-Llama3.1-8B 2.96 0.48 0.48 0.67
-ProvSyn 3.01 0.21 0.90 0.82
-Theia GPT-4o 2.39 0.65 0.89 0.53
-Claude-3.7 1.80 0.74 0.82 0.70
-Doubao-pro 2.78 1.43 0.56 0.39
-Qwen2.5-7B 1.36 8.90 0.74 0.48
-Llama3.1-8B 1.04 8.94 0.39 0.66
-ProvSyn 3.02 0.00 0.81 0.65
-Trace GPT-4o 2.16 0.46 0.70 0.27
-Claude-3.7 2.15 0.54 0.67 0.41
-Doubao-pro 2.09 0.40 0.57 0.08
-Qwen2.5-7B 2.10 0.47 0.75 0.09
-Llama3.1-8B 2.08 4.45 0.71 -0.01
-ProvSyn 2.80 0.10 0.81 0.75
-5.3 Semantic Correctness Evaluation
+Table 1 shows that across six datasets, ProvSyn achieves higher LCS scores and lower DTW values than the baselines in five datasets.
+In particular, for Cadets-E3, OpTC-H201, and OpTC-H501, ProvSyn attains LCS scores above 4.5, highlighting its effectiveness in maintaining temporal fidelity.
+Embedding Evaluation.
+For graph embedding, we use the netlsd library for NetLSD and Gensim’s Doc2Vec for Graph2Vec.
+It is important to note that absolute similarity values are generally unreliable due to variations in graph size, embedding scale, and initialization; instead, relative comparisons within the same dataset provide meaningful interpretation.
+As shown in Table 1, across six datasets, ProvSyn consistently achieves the highest relative similarity scores in five datasets for both NetLSD and Graph2Vec.
+Notably, on the OpTC datasets, where all baselines have NetLSD scores below 0.30, ProvSyn maintains scores above 0.60, demonstrating its robustness in preserving embedding fidelity.
+Semantic Evaluation In the experiment, prior to concatenating the BERT-derived node name embeddings with the GAT output embeddings, we employ a linear layer to project the BERT embeddings into the same dimensional space as the GAT outputs.
+This ensures prevents the concatenation process from being biased toward higher-dimensional components.
+Figure 3 shows that the provenance graphs generated by ProvSyn consistently outperform the baselines in semantic accuracy across all six datasets, demonstrating robustness to variations in data distribution.
+Notably, ProvSyn exhibits a substantial lead in the Theia-E3, ClearScope-E5, and OpTC-H501 datasets.
+For instance, on ClearScope-E5, ProvSyn achieves an accuracy of 0.98, significantly surpassing the best-performing baseline, Qwen3-max, which scores 0.74.
 Figure 3: Comparison of semantic accurary across different models and datasets. Higher values indicate better performance.
-We use the semantic correctness validation mechanism introduced in Section 4.6 to evaluate the semantic of synthetic graphs.
-In the experiment, the GAT model was trained for 2500 epochs with a learning rate of 0.0001, while the MLP classifier was trained for 1500 epochs with the same learning rate.
-To validate the effectiveness and interpretability of our semantic accuracy evaluation, we tested it on real-world provenance graphs from the same datasets but different from the training graphs.
-The results show that the semantic accuracy of these real graphs exceeds 90% across the Nodlink, Cadets, Theia, and Trace datasets.
-This supports the validity of our semantic accuracy evaluation method.
-The results in Figure 2 show that the provenance graphs generated by ProvSyn consistently outperform the baselines in semantic accuracy across all four datasets, showing robustness to differences in data distribution.
-In contrast, the baselines exhibit substantial fluctuations in semantic accuracy across datasets.
-Notably, in the more challenging Theia and Trace datasets, the semantic accuracy of baseline-generated graphs remained below 0.20, whereas ProvSyn achieved 0.57 and 0.82, respectively.
-These results demonstrate the robustness of ProvSyn in handling complex, low signal-to-noise provenance data.
-5.4 Data Imbalance Evaluation
-To evaluate whether the inclusion of ProvSyn-generated synthetic graphs mitigates class imbalance, we use two metrics—entropy and the Gini index—to quantify the distributional balance of node and edge labels.
-Given a label distribution with probabilities $p_{1},p_{2},\ldots,p_{n}$, the two metrics are defined as:
-$\text{Entropy}=-\sum_{i=1}^{n}p_{i}\log(p_{i}),\quad\text{Gini}=1-\sum_{i=1}^{n}p_{i}^{2}$
-Higher values of both metrics indicate a more uniform and balanced label distribution.
-We conduct experiments on the Nodlink, Cadets, Theia, and Trace datasets, both in their original data and after incorporating ProvSyn-generated data.
-For each dataset, we add 100 synthetic graphs merged with the original graph as graph communities.
-The results are summarized in Table V.
-As shown, the inclusion of ProvSyn increases both entropy and Gini index for node and edge labels across all datasets.
-In the Nodlink dataset, the metrics improve significantly due to the dataset’s small scale.
-These results suggest that ProvSyn contributes to a more balanced label distribution by enriching underrepresented classes in provenance graph datasets.
-To evaluate whether the synthesized graphs extend beyond the training data distribution and contain novel structures that cover more provenance scenarios, we use two metrics: Novelty and Uniqueness.
-Novelty is computed by verifying that each generated graph is neither a subgraph of any reference graph in the training set nor contains any reference graph as a subgraph, indicating whether novel structures exist in the synthesized graphs.
-Uniqueness measures the proportion of generated graphs that are structurally distinct from one another, reflecting the diversity of the generated graphs.
-As shown in Table VI, the graphs generated by ProvSyn achieve 100% novelty and 100% uniqueness across all four datasets.
+5.3 Data Imbalance Mitigation
+Table 3: Evaluation of data imbalance mitigation.
+The left section includes two metrics: Novelty (Nov.) and Uniqueness (Uniq.).
+The middle and right sections compare Node/Edge Normalized Entropy (Real vs.
+Real + Synthetic), where higher values indicate better balance.
+Dataset Nov. & Uniq. Node Entropy Edge Entropy
+Nov. Uniq. Real +ProvSyn Real +ProvSyn
+Cadets-E3 100.0 100.0 0.34 0.64 0.40 0.46
+Theia-E3 100.0 100.0 0.91 0.92 0.59 0.59
+ClearScope-E5 100.0 100.0 0.30 0.41 0.33 0.34
+Theia-E5 100.0 100.0 0.24 0.59 0.35 0.47
+OPTC-H201 100.0 100.0 0.43 0.70 0.56 0.57
+OPTC-H501 100.0 100.0 0.41 0.58 0.55 0.55
+Novelty and Uniqueness.
+As shown in Table 3, the graphs generated by ProvSyn achieve 100% novelty and 100% uniqueness across all four datasets.
 This indicates that the synthetic data exhibits strong diversity and effectively captures patterns that were underrepresented or missing in the original data.
-TABLE V: Comparison of node and edge label entropy and Gini index across different datasets with and without synthetic data. Higher values indicate better performance.
-Dataset Data Source Node Label Edge Label
-Entropy Gini Index Entropy Gini Index
-Nodlink Real 1.03 0.42 2.41 0.75
-+ProvSyn 1.31 (+0.28) 0.55 (+0.13) 2.63 (+0.22) 0.82 (+0.07)
-Cadets Real 1.11 0.44 1.84 0.52
-+ProvSyn 1.31 (+0.20) 0.53 (+0.09) 1.92 (+0.08) 0.54 (+0.02)
-Theia Real 1.31 0.53 1.33 0.39
-+ProvSyn 1.41 (+0.10) 0.58 (+0.05) 1.36 (+0.03) 0.39 (+0.00)
-Trace Real 1.52 0.57 1.57 0.45
-+ProvSyn 1.62 (+0.10) 0.61 (+0.04) 1.61 (+0.04) 0.47 (+0.02)
-TABLE VI: Evaluation of ProvSyn-generated graphs in terms of Novelty and Uniqueness across different datasets. Higher values indicate better performance.
-Dataset Novelty (%) Uniqueness (%)
-Nodlink 100.0 100.0
-Cadets 100.0 100.0
-Theia 100.0 100.0
-Trace 100.0 100.0
+Normalized Entropy.
+We compute normalized entropy for both the original datasets and the augmented datasets after incorporating ProvSyn-generated graphs.
+The results are summarized in Table 3.
+Overall, the inclusion of ProvSyn-generated data consistently increases the normalized entropy of node and edge label distributions across all datasets.
+Notably, datasets with more severe initial imbalance exhibit more pronounced entropy improvements after augmentation.
+For example, in Theia-E5, the node label normalized entropy increases substantially from 0.24 to 0.59.
+These results suggest that ProvSyn contributes to a more balanced label distribution by enriching underrepresented classes.
 5.4 Application in IDS
-To demonstrate the practical utility of ProvSyn, we focus on the task of APT attack detection.
-We conduct experiments using two detection algorithms, Magic[31] and Nodlink[37].
-Specifically, we augment the training set with synthesized provenance graphs and then observe whether the detection performance improves or declines during testing.
-For the Nodlink algorithm, we enhanced its dataset by adding 50 synthetic graphs merged with the original provenance graph as graph communities.
-This resulted in a 15% increase in node counts.
-For the Magic algorithm, we enhanced the Cadets, Theia, and Trace datasets.
-Due to their larger scale, we added 300 synthetic graphs merged into their original provenance graphs.
-This led to a 10% to 15% increase in node counts.
-To evaluate the effectiveness of synthetic graphs in intrusion detection tasks, we maintain identical training configurations before and after incorporating synthetic data.
-The detection performance is presented in Table VII.
-In addition to Precision and Recall, we report the number of False Positives (#FP), an crucial metric for intrusion detection system, as a high false positive rate can increase the workload of subsequent security analysis.
-As shown in the results, on the Nodlink dataset, the inclusion of synthetic graphs leads to no change in Precision, Recall, or the number of false positives, indicating that the synthetic data is of comparable quality to the original data.
-On the Cadets, Theia, and Trace datasets, Recall remains constant, and Precision improves.
-Specifically, false positives decrease by 506 on Cadets (from 6435 to 5929), by 95 on Theia (from 931 to 836), and by 379 on Trace (from 2390 to 2011).
-These results suggest that augmenting with synthetic graphs produced by ProvSyn improves model performance, and effectively reduces false positives in detection systems.
-TABLE VII: IDS performance across different datasets with and without synthetic data. Higher values indicate better performance.
-Dataset Data Source Precision(%) Recall(%) #FP
-Nodlink Real 100.00 100.00 0
-+ ProvSyn 100.00 100.00 0
-Cadets Real 66.58 99.79 6435
-+ ProvSyn 68.37 99.77 5929
-Trace Real 96.61 99.98 2390
-+ ProvSyn 97.13 99.98 2011
-Theia Real 96.45 99.99 931
-+ ProvSyn 96.80 99.99 836
+Table 4: Performance comparison of IDS models trained on different data compositions.
+The “Real” rows represent models trained solely on real data, while “+ ProvSyn” rows represent models trained on a combination of real and synthesized provenance data.
+Green cells indicate performance improvement over the baseline, and Red cells indicate a decrease in performance.
+(F-Score and MCC are scaled by $10^{3}$)
+Nodlink Flash
+Dataset Source F-Score B. Acc MCC DOR F-Score B. Acc MCC DOR
+Cadets-E3 Real 1.72 0.75 19.83 9.74 0.00 0.50 0.00 0
++ ProvSyn 1.94 0.84 25.41 35.24 4.85 0.59 20.47 12.55
+Theia-E3 Real 0.11 0.47 -2.55 0.31 0.22 0.44 -3.24 0.56
++ ProvSyn 0.60 0.53 2.71 1.89 0.44 0.53 1.99 1.41
+Clearscope-E5 Real 0.29 0.47 -3.27 0.41 0.58 0.48 -1.44 0.82
++ ProvSyn 0.97 0.61 7.93 2.50 0.83 0.57 5.53 2.00
+Theia-E5 Real 1.16 0.73 15.56 12.51 0.13 0.35 -14.90 0.08
++ ProvSyn 1.03 0.80 17.20 18.63 0.30 0.60 4.08 2.29
+OpTC-H201 Real 32.24 0.72 671.63 15.61 0.11 0.17 N/A 0.01
++ ProvSyn 32.61 0.72 1446.15 15.81 5.12 0.55 293.88 1.53
+OpTC-H501 Real 6.10 0.70 N/A 10.73 0.15 0.26 -335.60 0.07
++ ProvSyn 6.82 0.72 225.50 12.89 0.86 0.46 -35.13 0.73
+Magic Velox
+Dataset Source F-Score B. Acc MCC DOR F-Score B. Acc MCC DOR
+Cadets-E3 Real 0.99 0.73 14.61 12.92 0.00 0.50 -0.04 0
++ ProvSyn 1.54 0.80 20.76 17.49 3.48 0.51 6.43 7.88
+Theia-E3 Real 0.42 0.57 N/A 2.00 54.61 0.57 67.78 243.04
++ ProvSyn 0.63 0.72 N/A 12.46 140.35 0.55 151.65 1934.86
+Clearscope-E5 Real 0.63 0.48 -1.20 0.88 48.19 0.52 49.25 205.28
++ ProvSyn 0.86 0.61 9.32 15.30 174.76 0.59 174.48 751.84
+Theia-E5 Real 0.30 0.69 7.45 44.23 54.79 0.51 120.37 11192.40
++ ProvSyn 0.39 0.76 10.07 79.14 56.34 0.51 170.24 $\infty$
+OpTC-H201 Real 1.67 0.27 -1882.50 0.13 0.68 0.50 3.61 21.48
++ ProvSyn 1.67 0.28 -915.75 0.15 0.69 0.50 4.82 38.01
+OpTC-H501 Real 0.42 0.26 N/A 0.11 2.65 0.50 13.77 331.65
++ ProvSyn 2.04 0.63 N/A 3.04 2.66 0.50 21.07 994.95
+To demonstrate the practical utility of ProvSyn, we focus on APT attack detection using four representative systems: Nodlink [li2023nodlink], Flash [rehman2024flash], Magic [jia2024magic], and Velox [bilot2025sometimes].
+All systems are reproduced using the open-source PIDSMaker [bilot2026pidsmaker, bilot2025sometimes] framework.
+Our evaluation focuses on whether augmenting the training set with synthesized provenance graphs enhances detection performance.
+To this end, we incorporate synthetic graphs, as graph communities, into every time window of the training data from the first day.
+The detection performance is presented in Table 4.
+Based on the results, we have the following four key findings.
+Detection Performance Enhancement.
+The integration of ProvSyn data consistently strengthens detection metrics across various baselines, demonstrating its efficacy in both high-performing and low-performing scenarios.
+For instance, when a model already exhibits reasonable performance, such as Nodlink on the Cadets-E3 dataset, the B.
+Acc increases from $0.75$ to $0.84$ and the F-Score improves from $1.72\times 10^{-3}$ to $1.94\times 10^{-3}$.
+More importantly, the augmentation provides a critical correction in cases where the initial performance is near or below random chance.
+A prominent example is observed with the Flash algorithm on the OpTC-H201 dataset, where the B.
+Acc rises significantly from $0.17$ to $0.55$.
+Similar improvements are seen in the OpTC-H501 dataset for Flash, where the B.
+Acc progresses from $0.26$ to $0.46$, effectively rescuing the model from a state of total detection failure.
+Discriminative Reliability and Diagnostic Utility.
+The impact of ProvSyn is further evidenced by the improvement in MCC and DOR, which reflect the model’s correlation with the ground truth and its diagnostic strength.
+The MCC values are consistently pulled toward positive territory, particularly in instances where the baseline models initially yield negative results, such as Nodlink on Theia-E3 moving from 2.55-2.55 to $2.71$ and Flash on Clearscope-E5 shifting from 1.44-1.44 to $5.53$.
+This indicates a fundamental shift from biased or inverse classification to a correct positive correlation.
+Simultaneously, the DOR increases across all experiments, often by several orders of magnitude.
+For example, the DOR for Velox on Theia-E3 jumps from $243.04$ to $1934.86$, and on Theia-E5, it reaches $\infty$.
+These results suggest that the diagnostic efficiency for identifying attack samples is substantially sharpened through the augmented training set.
+Multi-Factor Sensitivity Analysis.
+The efficacy of ProvSyn is closely coupled with three key factors.
+Regarding the IDS architecture, the improvement for Flash on OpTC-H201 is more pronounced than for Nodlink, which may be because the baseline detection performance of Flash is lower than that of Nodlink, providing a larger margin for potential enhancement.
+For dataset characteristics, while Magic shows negligible F-Score improvement on OpTC-H201, it demonstrates significant F-score improvements across the other five datasets.
+Finally, metric-specific sensitivity is observed in Velox; while its B.Acc on OpTC-H201 remains stable at $0.50$, its MCC and DOR show notable increases from $3.61$ to $4.82$ and $21.48$ to $38.01$ respectively, suggesting that while the absolute detection threshold may be constrained by the algorithm’s architecture, its discriminative confidence and diagnostic power are still bolstered by the augmented data.
+Enhancement Mechanism Analysis.
+The overall performance uplift across different dataset–IDS combinations can be attributed to the mitigation of data imbalance and the diversity introduced by ProvSyn synthesized samples.
+By incorporating a wider variety of benign provenance graphs into the training set, the models can cover previously underrepresented system behaviors and edge-case scenarios.
+This expanded representation of the benign space helps the detection algorithms to more accurately define the boundaries of normal activity.
+Consequently, the models become less prone to overfitting on limited real-world benign data, enabling them to better distinguish subtle malicious deviations from legitimate but rare system actions, thereby improving the robustness of APT detection.
 5.5 Overhead Analysis
-TABLE VIII: Time and memory requirements of ProvSyn’s core modules.
-Module Phase Time Memory
-Graph Generation Training 18 hours 0.84 GB
-Inference 28.09 s 0.38 GB
-Text Generation Training 109 s 3.44 GB
-Inference 2.78 hours 28 GB
-In this section, we analyze the time and memory requirements of ProvSyn (see Table VIII).
-The most time- and resource-intensive components of the ProvSyn framework are the heterogeneous graph generation module and the text attribute generation module.
-For the heterogeneous graph generation module, we use 16,000 samples for training, with a batch size of 32 and 3,000 training epochs.
+We analyze the time and memory requirements of ProvSyn.
+The most time- and resource-intensive components of the ProvSyn framework are the heterogeneous graph synthesis module and the text attribute synthesis module.
+For the graph structure synthesis module, we use 16,000 samples for training, with a batch size of 32 and 3,000 training epochs.
 The average training time is 18 hours, with a GPU memory usage of 0.84 GB.
 Inference for generating 1,000 graphs takes 28.09 seconds and consumes 0.38 GB of memory.
-For the text attribute generation module, we adopt Llama3.2-3B-Instruct as the base model, applying the LoRA algorithm with 4-bit quantization.
+For the text attribute synthesis module, we adopt Llama3.2-3B-Instruct as the base model, applying the LoRA algorithm with 4-bit quantization.
 The FastLanguageModel library is used to accelerate training.
@@
 This analysis demonstrates that ProvSyn is a time- and memory-efficient framework for provenance graph synthesis.
+6 Related Work
+Provenance-based Intrusion Detection System (pids).
+Provenance-based detection methods can be divided into three main categories [zengy2022shadewatcher]: rule-based, statistics-based, and learning-based techniques.
+Rule-based methods [hassan2020tactical, milajerdi2019holmes, hossain2020combating] rely on predefined heuristics derived from known attack patterns to identify malicious activities.
+Statistics-based [hassan2019nodoze, wang2020you, liu2018towards] approaches analyze deviations in provenance graph elements by measuring their statistical anomalies.
+Learning-based techniques employ deep learning models to capture either normal system behavior [han2020unicorn, wang2022threatrace] or malicious patterns [liu2019log2vec, zengy2022shadewatcher, li2021hierarchical], framing APT detection as either a classification [li2021hierarchical] or anomaly detection task [zengy2022shadewatcher, han2021sigl].
+These learning-based methods further branch into sequence-based approaches [liu2019log2vec], which focus on execution workflows, and graph-based [han2020unicorn, wang2022threatrace, zengy2022shadewatcher, han2021sigl, li2021hierarchical] methods, which use graph neural networks to model entity relationships and detect behavioral anomalies.
+LLM-driven Synthetic Data Generation.
+LLM-based synthetic data generation methods can be categorized into three main types: prompt engineering-based approaches [gilardi2023chatgpt, gunasekar2023textbooks, sudalairaj2024lab, liu2024best, yoo2021gpt3mix, li2023textbooks, ye2022progen], which utilize task descriptions, condition-value prompts, and in-context examples to steer generation; multi-step generation-based approaches [wei2022chain, gao2023retrieval, ding2023enhancing, wang2022self, honovich2022unnatural], which break down complex tasks either at the sample or dataset level; and knowledge-enhanced approaches [pan2024unifying, chen2024mindsearch, ji2021survey], which incorporate external knowledge such as knowledge graphs or web sources to improve the factual quality of outputs.
+The evaluation of synthetic data includes both direct metrics, which measure faithfulness [lee2022factuality] and diversity [yu2023large], and indirect metrics, which assess performance on downstream tasks through benchmark testing [sun2023principle] and open-ended evaluation using human [he2023annollm] or model-based [xu2023wizardlm] methods.
 7 Discussion and Limitations
@@
 A provenance graph is not only a directed heterogeneous graph with textual attributes, but also a temporal graph in which each event is associated with a timestamp indicating its occurrence time in the system.
-These temporal information has recently been leveraged by downstream intrusion detection systems [53] to better capture the patterns of benign and anomalous behaviors.
+These temporal information has recently been leveraged by downstream intrusion detection systems [rehman2024flash] to better capture the patterns of benign and anomalous behaviors.
 However, in ProvSyn, we do not synthesize explicit timestamps for events.
 Instead, we approximate temporal order by relying on the sequence of directed edges.
-Future work may explore modeling temporal dynamics using temporal graph generation networks [75], or designing strategies that leverage LLMs to generate realistic timestamps, thereby enabling more comprehensive and temporally coherent provenance graph synthesis.
+Future work may explore modeling temporal dynamics using temporal graph generation networks [zhang2020tg], or designing strategies that leverage LLMs to generate realistic timestamps, thereby enabling more comprehensive and temporally coherent provenance graph synthesis.
 Graph Scale.
@@
 To enable data augmentation, we embed multiple synthetic graphs as distinct communities within large real-world provenance graphs.
-The results show that this approach effectively mitigates data imbalance and reduces the false positive rate of detection models trained on the augmented datasets.
+The results show that this approach effectively mitigates data imbalance and improves accuracy of detection models trained on the augmented datasets.
 Future work can focus on scaling up synthetic provenance graphs by enhancing graph generation models for large-scale synthesis and exploring graph merging algorithms that combine smaller graphs into larger ones while preserving structural fidelity.
-Efficiency of Text Attribute Synthesis.
-During textual attribute synthesis, overlaps among nodes in different sequences generated by DFS can lead to naming conflicts when these sequences are processed in parallel to generate node names.
-To address this issue, we serialize the sequences and input them into the LLM sequentially.
-The generated node names from earlier sequences are then used as context for subsequent sequences.
-This approach is effective for small-scale graphs.
-However, the synthesis efficiency decreases for large-scale graphs.
-Future work may explore strategies to partition large graphs into batches of non-overlapping sequences, enabling parallel processing while avoiding node naming conflicts.
 Application Scenarios.
@@
 8 Conclusion
-This paper proposes ProvSyn, a novel provenance graph generation framework consisting of three stages: heterogeneous graph structure generation, rule-based topological refinement, and LLM-based synthesis of textual attributes.
-To evaluate the fidelity of the generated graphs, we assess structural, textual, temporal, embedding, and semantic aspects.
-Compared with provenance data directly synthesized by LLMs, ProvSyn consistently produces graphs with higher fidelity.
-When integrated into existing datasets, ProvSyn-generated graphs help mitigate class imbalance and lead to reduced false positive rates in detection models trained on the augmented data.
-Appendix A Entity and Event Distribution in Our Datasets
-We present the distributions of entity type and event type across four commonly-used public provenance datasets in Figure 3.
+This paper proposes ProvSyn, a novel provenance graph generation framework consisting of three compoments: graph structure synthesis via heterogeneous graph generation models, textual attribute synthesis via fine-tuned LLM, and five-dimensional fidelity evaluation.
+Compared to provenance data synthesized directly by four state-of-the-art LLMs, ProvSyn consistently produces graphs with higher fidelity.
+When integrated into APT detection training sets, ProvSyn-generated graphs effectively mitigate class imbalance and yield improved accuracy.
+Ethical Considerations
+To the best of our knowledge, this work does not present any ethical concerns.
+All experiments were conducted on publicly available datasets that were obtained in compliance with ethical standards and do not contain any sensitive or personally identifiable information.
+The use of large language model APIs in this research adhered strictly to the official terms of service and usage policies provided by the respective providers.
+Open Science
+To to facilitate future research in provenance graph synthesis and pids, we release the source code of ProvSyn at https://anonymous.4open.science/r/OpenProvSyn-4D0D/.
+Appendix A Entity and Event Distribution
+We present the distributions of entity type and event type across six commonly-used public provenance datasets in Figure 4.
 The distributions exhibit a typical long-tail pattern and reveal significant data imbalance.
-Appendix B Textual Quality Evaluation
-In Algorithm 2, we compute the maximum similarity score between the text attribute of each node and the reference corpus.
-This metric serves to evaluate the model’s maximum potential in generating accurate and relevant attributes.
-Input : Generated graph nodes $V_{\text{gen}}=\{g_{1},g_{2},...,g_{m}\}$ with predicted names and types;
-Real graph nodes $V_{\text{real}}=\{r_{1},r_{2},...,r_{n}\}$ with ground-truth names and types;
-Node type set $\mathcal{T}=\{t_{1},t_{2},...,t_{k}\}$
-Output : Final textual quality score $S_{\text{final}}$
-1
-2foreach $t\in\mathcal{T}$ do
-       $\mathcal{R}_{t}\leftarrow\{r_{j}\mid r_{j}\in V_{\text{real}},\text{type}(r_{j})=t\}$
-        // Construct reference corpus
-3
-[…]
-[…]
-6
-7foreach $g_{i}\in V_{\text{gen}}$ do
-8       $t\leftarrow\text{type}(g_{i})$
-[…]
-10       foreach $r_{j}\in\mathcal{R}_{t}$ do
-[…]
-[…]
-[…]
-14
-[…]
-[…]
-17
-[…]
-  // Aggregate final score
-18
-19 return $S_{\text{final}}$
-20
-Algorithm 2 Textual Quality Evaluation
-Appendix C Hyperparameters Setting
-Figure 3: Entity Type and Event Type Distribution in Provenance Dataset including Nodlink, Cadets, Theia and Trace.
-Figure 4: Hidden size setting in Nodlink dataset.
-Figure 5: Inference temperature setting in Nodlink dataset.
-Figure 6: Ablation Study.
-Hidden size is of LSTM for heterogeneous graph generation.
-Hidden size is a key parameter for LSTM, as it determines the capacity of the model to capture temporal and structural dependencies in sequential node and edge generation.
-We conducted experiments on the Nodlink dataset and evaluated structural fidelity using five MMD-based metrics in the validation set.
-The results in Figure 4 show that a hidden size of 64 yields significantly lower fidelity compared to the other settings.
+Appendix B Post-processing Rules
+Table 5: Dataset-specific post-processing schemas
+Dataset Node Pair (Src, Dst) Permissible Event Types
+DARPA E3/E5 (subject, subject) read, write, open, connect, recvfrom, sendto, clone, sendmsg, recvmsg
+(subject, file) write, connect, sendmsg, sendto, clone
+(subject, netflow) write, sendto, connect, sendmsg
+(file, subject) read, open, recvfrom, execute, recvmsg
+(netflow, subject) open, read, recvfrom, recvmsg
+OpTC (subject, subject) create, open, terminate
+(subject, file) create, delete, modify, rename, write
+(file, subject) read
+(subject, netflow) message, start
+(netflow, subject) message, open, start
+For each dataset, we define a dataset-specific post-processing schema that constrains the permissible event types between different pairs of node types, as summarized in Table 5.
+These schemas are derived from the structural regularities observed in the corresponding ground-truth provenance graphs.
+Appendix C Baseline LLM Prompt
+We present in Figure 7 the prompt used to instruct LLMs in generating provenance logs on Cadets-E3 dataset.
+In our experiments, the Examples section contains 20 examples.
+For presentation brevity, only 5 examples are shown.
+Appendix D Hyperparameters Setting
+Hidden size is of LSTM for graph structure synthesis.
+Hidden size is a key parameter for LSTM, as it determines the capacity of the model to capture structural dependencies in sequential node and edge generation.
+We conducted experiments on the Theia-E3 dataset and evaluated structural fidelity using five MMD-based metrics in the validation set.
+The results in Figure 5 (a) show that a hidden size of 64 yields significantly lower fidelity compared to the other settings.
 For hidden sizes of 128, 256, and 512, the MMD values remain at a comparable level, with 256 performing slightly better than 128 and 512.
@@
 Higher temperatures yield more diverse outputs, while lower temperatures produce more deterministic results.
-We conducted experiments on the Nodlink dataset using temperature values of 0.1, 0.5, 1.0, and 1.5, with all other inference parameters held constant.
-Evaluation focused on the quality of text generated for process and file node attributes, as well as semantic correctness.
-As shown in Figure 5, the best performance is achieved when the temperature is set to 1.5.
-In particular, at lower temperatures, the generated node names tended to be repetitive and lacked diversity.
-Based on these findings, we selected 1.5 as the final inference temperature for ProvSyn, consistent with the recommendation of the FastLanguageModel library.
-Appendix D Ablation Study
+We conducted experiments on the Cadets-E3 dataset using temperature values of 0.1, 0.5, 1.0, and 1.5, with all other inference parameters held constant.
+Evaluation focused on the quality of text generated for process and file node attributes.
+As shown in Figure 5 (b) and (c), the best performance is achieved when the temperature is set to 1.5.
+Based on these findings, we selected 1.5 as the final inference temperature for ProvSyn.
+Appendix E Ablation Study
 Alternative models to GraphGen.
-In the heterogeneous graph generation module, we adopt the design of GraphGen, which encodes graphs as DFS code sequences and employs an LSTM for training and inference.
-To validate the effectiveness of this design, we compare it with alternative models, including the widely used GraphRNN [72].
-In our experiments, we evaluate the graphs generated by GraphRNN and GraphGen using five MMD metrics.
-As shown in Figure 6(a), the graphs generated by GraphGen consistently achieve significantly lower MMD scores across all five metrics, indicating substantially better fidelity compared to those generated by GraphRNN.
+For heterogeneous graph structure generation, we adopt the design of GraphGen, which encodes graphs as DFS code sequences and employs an LSTM for training and inference.
+To validate the effectiveness of this design, we compare it with alternative models, including the widely used GraphRNN [you2018graphrnn].
+In our experiments, we evaluate the graphs generated by GraphRNN and GraphGen using four MMD metrics.
+As shown in Figure 6(a), the graphs generated by GraphGen consistently achieve lower MMD scores across all MMD metrics, indicating better fidelity compared to those generated by GraphRNN.
 These results demonstrate the effectiveness of using GraphGen for heterogeneous graph generation in ProvSyn.
@@
 To evaluate the effectiveness of the masking design, we perform ablation experiments by disabling one of the strategies and using only the other.
-Model performance is then evaluated using text quality metrics and semantic correctness.
-The experiments are conducted on the Trace dataset, with the number of training samples fixed at 20,000 for both the single-strategy and dual-strategy settings.
+Model performance is then evaluated using textual and semantic metrics.
+The experiments are conducted on the Clearscope-E5 dataset, with the number of training samples fixed at 20,000 for both the single-strategy and dual-strategy settings.
 In the results shown in Figure 6(b), we observe that using only the Full Masking strategy yields better performance than using only the Part Masking strategy.
 This may be because the Part Masking strategy provides partial node names within the sequence, offering more cues for the LLM to leverage.
-As a result, a model trained solely with Part Masking performs poorly when tested on Full Masking sequences, which contain fewer clues and less information.
-In contrast, using both strategies together achieves better performance than either alone, as it exposes the model to a broader range of sequence types that may appear during inference.
+As a result, a model trained solely with Part Masking performs poorly when tested on Full Masking sequences, which contain fewer clues.
+In contrast, using both strategies together achieves better performance than either alone, as it exposes the model to a wider range of sequence types that may appear during inference.
 These results demonstrate the effectiveness of our proposed approach that combines both masking strategies.
@@
 We also compare its performance against models of different sizes, including Llama-3.2-1B-Instruct and Llama-3.1-8B-Instruct.
-All models are trained on the same training datasets and evaluated using both text quality metrics and the semantic correctness metric.
-The results, presented in Figure 6(c), demonstrate that the 3B model achieves the highest scores on the BLEU and GLEU metrics.
-In contrast, the 8B model shows marginally better performance on the ROUGE and semantic correctness metrics.
+All models are trained on the same training datasets and evaluated using textual and semantic metrics.
+The results, presented in Figure 6(c), demonstrate that the 3B model achieves the highest scores on the Distinct-1 and semantic accuracy.
+In contrast, the 8B model shows marginally better performance on the GLEU.
 Given that the advantage of the 8B model is not substantial and considering computational resource constraints, we select the 3B model as the base for synthesizing text attributes.
 In practice, ProvSyn introduces a general training strategy to leverage LLMs for synthesizing text attributes, where the choice of the base model involves balancing the complexity of the dataset, the application requirements and the limitations of computational resources.
+Figure 4: Entity Type and Event Type Distribution in Provenance Dataset.
+Figure 5: Hidden size setting and inference temperature setting.
+Figure 6: Ablation Study.
+[Task Description] You are an audit log generator.
+Carefully analyze the audit log records provided in [Examples] to understand their structure, semantics, and temporal patterns.
+Based on this understanding, generate 50 unique audit log records that are logically consistent, structurally valid, and realistic.
+The generated logs should reflect plausible system activities, preserve coherent relationships between entities and events, and follow realistic ordering observed in the examples.
+[Examples]
+"event_type": "EVENT_OPEN",
+"source_node_type": "file",
+"source_node_name": "kernel_mem.txt",
+"target_node_type": "subject",
+"target_node_name": "vmstat"
+"event_type": "EVENT_READ",
+"source_node_type": "file",
+"source_node_name": "None",
+"target_node_type": "subject",
+"target_node_name": "pickup"
+"event_type": "EVENT_WRITE",
+"source_node_type": "subject",
+"source_node_name": "head",
+"target_node_type": "file",
+"target_node_name": "top_procs.txt"
+"event_type": "EVENT_CONNECT",
+"source_node_type": "subject",
+"source_node_name": "master",
+"target_node_type": "file",
+"target_node_name": "None"
+"event_type": "EVENT_EXECUTE",
+"source_node_type": "file",
+"source_node_name": "/usr/bin/top",
+"target_node_type": "subject",
+"target_node_name": "top"
+[Output] Return the logs in JSON format. Only output the JSON.
+Figure 7: Prompt for LLM Provenance Data Generation on Cadets-E3